### v2

# Rui Lia,  Anyao Wangb   Mingqing Zhaic

Abstract
A graph $G$ is said to satisfy the Hereditarily Bounded Property $P_{t,r}$ if, for every subgraph $H$ of $G$ with at least $t$ vertices, the size of $H$ satisfies $|E(H)|\leq t\cdot|V(H)|+r$. In 1996, Guiduli posed the following problem: If an $n$-vertex graph $G$ possesses the property $P_{t,r}$, where $t$ is a positive real number and $r$ is a real number such that $r\geq-\binom{\lfloor t+1\rfloor}{2}$, does it follow that the spectral radius $\rho(G)$ is bounded above by $c(s,t)+\sqrt{\lfloor t\rfloor n}$, where $s=\binom{t+1}{2}+r$?
In this paper, we affirmatively resolve Guiduli’s problem. Furthermore, we provide a precise characterization of the extremal graphs that achieve this bound. Specifically, these extremal graphs are constructed as the join of two graphs: one being a complete graph $K_{\lfloor t\rfloor}$, and the other either a disjoint union of $K_{3}$ and isolated vertices, or a forest consisting solely of star structures.
The forest of order $n-\lfloor t\rfloor$ in extremal graphs is meticulously characterized. The key to our proof is the classification of the stars into four categories–maximal, large, small, and minimal–according to their sizes. This classification provides deeper insights into how edge distribution influence the spectral properties of graphs under the constraint of property $P_{t,r}$.
Keywords: Hereditarily Bounded Property; spectral radius; extremal graph
AMS   Classification: 05C50; 05C35
1 Introduction
All graphs considered in this context are simple, undirected and free from isolated vertices. We use $|G|$ and $e(G)$ to denote the numbers of vertices and edges in a graph $G$, respectively. For a subset $S$ of $V(G)$, let $G[S]$ be the subgraph induced by $S$. Let $N_{U}(v)$ denote the set of neighbors of vertex $v$ within the subset $U$ of vertices, formally defined as $N_{U}(v)=\{u\mid u\in U\text{ and }uv\in E(G)\}$. In particular, when $U=V(G)$, we simply write $N(v)$ for $N_{V(G)}(v)$. Denoted by $d_{U}(v)$ the cardinality of the set $N_{U}(v)$. Moreover, let $e(U)$ denote the size of $G[U]$ and $e(U,W)$ be the number of edges connecting $U$ and $W$ in $G$. Let $kG$ be the disjoint union of $k$ copies of $G$, and $G_{1}\nabla G_{2}$ be the graph obtained from $G_{1}$ and $G_{2}$ by joining each vertex of $G_{1}$ with each vertex of $G_{2}$.
Let $A(G)$ denote the adjacency matrix of a graph $G$, and let $\rho(G)$ represent the spectral radius of $A(G)$. A graph $G$ is said to have the Hereditarily Bounded Property $P_{t,r}$ if for each subgraph $H\subseteq G$ with $|V(H)|\geq t$, the size of $H$ satisfies $|E(H)|\leq t\cdot|V(H)|+r$, where $t>0$. In fact, many classes of graphs possess the Hereditarily Bounded Property, including planar graphs, which have been extensively studied in the context of spectral extremal problems (see [1, 3, 6, 7, 8, 15, 20]). Research into the spectral radius of graphs with similar hereditary properties has also flourished. For instance, constraints on matching numbers have been explored by Feng et al. [11], while constraints on cycle lengths were investigated by Gao et al. [12]. These studies contribute to a broader understanding of how the structural properties of graphs influence their spectral characteristics.
This concept was first introduced by Guiduli in his Ph.D. thesis [13], wherein he proved the following theorem.
Theorem 1.1.
Let $t\in\mathbf{N}$ and $r\geq-\binom{t+1}{2}$. If $G$ is a graph on $n$ vertices with the property $P_{t,r}$, then $\rho(G)\leq(t-1)/2+\sqrt{t(t+1)+2r}+\sqrt{tn}$.
For ease of notation, Guiduli defined $s=\binom{t+1}{2}+r$ in his thesis. When $t$ and $r$ are not restricted to integers, this result implies that $\rho(G)\leq c(s,t)+\sqrt{\lceil t\rceil n}$, where $c(s,t)$ represents a function of $s$ and $t$. Guiduli also posed the following problem.
Problem 1.1.
Let $t,r\in\mathbf{R}$, $t>0$ and $r\geq-\binom{\lfloor t+1\rfloor}{2}$. Is it true that $G$ has the property $P_{t,r}$, then $\rho(G)\leq c(s,t)+\sqrt{tn}$?
Furthermore, Guiduli conjectured a potentially stronger statement:
$\rho(G)\leq c(s,t)+\sqrt{\lfloor t\rfloor n}.$
In this paper, we affirmatively resolve the conjecture posed by Guiduli. Specifically, we establish the following theorem:
Theorem 1.2.
Let $t,r\in\mathbf{R}$, $t>0$ and $r\geq-\binom{\lfloor t+1\rfloor}{2}$. If $G$ is an $n$-vertex graph with the property $P_{t,r}$ , then $\rho(G)\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$.
Finally, we consider the structure of $G$ when it attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. The following theorem provides a precise characterization of such graphs.
Theorem 1.3.
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. Then $G$ is the join of two graphs: one is $K_{\lfloor t\rfloor}$, another is either $K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$ or a forest consisting of stars.
In Section 4, we provide a detailed characterization of the forest structure, classifying the stars into maximal, big, small, and minimal stars based on their sizes.
2 Stability theorem
Let $\mathcal{G}$ denote the family of $n$-vertex graphs that satisfy property $P_{t,r}$. For any $G\in\mathcal{G}$, we assume for simplicity that the graph $G$ considered in this paper is connected. If $G$ were not connected, we could identify a component $C$ within $G$ such that the induced subgraph $G[C]$ also satisfies property $P_{t,r}$ and has the same spectral radius as $G$, i.e., $\rho(G[C])=\rho(G)$. Therefore, it suffices to consider only the connected component $G[C]$ for our analysis. Furthermore, we may assume that $n$ is sufficiently large. If $n$ were a finite constant, we could set $c(s,t)=n$. In this case, the spectral radius would be bounded by $\rho(G)\leq\rho(K_{n})=n-1\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$. Consequently, Theorem 1.2 would hold trivially.
Thus, for the remainder of our discussion, we focus on connected graphs $G\in\mathcal{G}$ with a sufficiently large order $n$. Given that $G\in\mathcal{G}$, it implies that $t\geq 1$. To see why, suppose for contradiction that $10<t<$ 1. For sufficiently large $n$, $e(G)\leq tn+r<n-1$. However, since $G$ is connected, it must have at least $n-1$ edges. This leads to a contradiction, hence $t\geq 1$.
By the well-known Perron-Frobenius theorem, there exists a positive unit eigenvector $X$ corresponding to the spectral radius $\rho(G)$. Let $X=(x_{1},x_{2},\ldots,x_{n})^{T}$ and let $u^{*}\in V(G)$ such that $x_{u^{*}}=\max_{u\in V(G)}x_{u}.$ For simplicity, we will write $\rho=\rho(G)$.
For convenience, we define $s=\big{\lceil}\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\big{\rceil}+1$ in this paper. This definition of $s$ differs from that given in Guiduli’s thesis but similarly depends on $t$ and $r$. It is clear that $s>t+1$ as $r\geq-\binom{\lfloor t+1\rfloor}{2}$. To prove our main result, we introduce a stability theorem that plays a crucial role in our analysis.
Theorem 2.1.
Let $G$ be a graph of sufficiently large order $n$. Let $X$ be a non-negative eigenvector corresponding to $\rho(G)$, and $x_{u^{*}}=\max_{u\in V(G)}x_{u}$. If $G$ has property $P_{t,r}$ and $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, then there exists a vertex subset $V$ of $G$ with $|V|=\lfloor t\rfloor$ such that $x_{v}\geq(1-\frac{1}{50s^{2}})x_{u^{*}}$ and $d_{G}(v)\geq(1-\frac{1}{25s^{2}})n$ for every $v\in V$.
Next, we will prove Theorem 2.1. We’ll define the set of vertices with large entries in $X$ relative to $x_{u^{*}}$, provide an upper bound for its cardinality, and then determine its exact value. This method was first introduced by Cioabă et al. to address a conjecture by Nikiforov [18], which comprises two parts: a spectral version of Erdős-Sós conjecture [4] and the spectral even cycle problem [5]. It was further developed by Byrne et al. [2].
Let $V_{1}=\{v\in V(G):\ x_{v}\geq(5s)^{-4}x_{u^{*}}\}$, $V_{2}=\{v\in V(G):\ x_{v}\geq(5s)^{-5}x_{u^{*}}\}$ and $V=\{v\in V(G):\ x_{v}\geq(5s)^{-1}x_{u^{*}}\}$. Obviously, $V\subseteq V_{1}\subseteq V_{2}$.
Claim 2.1.
$|V_{1}|\leq(5s)^{-6}n$ and $|V_{2}|\leq(5s)^{-5}n$.
Proof.
Let $v\in V_{1}$. We start with the inequality
$\rho x_{v}\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}(5s)^{-4}x_{u^{*}}.$
Moreover, given that $n$ is sufficiently large, it follows that $\rho x_{v}\geq 2s(5s)^{6}x_{u^{*}}.$ Additionally, we have $\rho x_{v}=\sum_{u\in N(v)}x_{u}\leq|N(v)|x_{u^{*}},$ where $N(v)$ denotes the set of neighbors of $v$ in $G$. Therefore,
$\sum_{v\in V_{1}}\rho x_{v}\leq\sum_{v\in V_{1}}|N(v)|x_{u^{*}}\leq\sum_{v\in V(G)}|N(v)|x_{u^{*}}=2e(G)x_{u^{*}}.$
Since $G\in\mathcal{G}$ and $s>t+1$, we know that $e(G)\leq tn+r\leq sn$. Combining this with the previous inequalities, we obtain
$2s(5s)^{6}|V_{1}|\leq\sum_{v\in V_{1}}\rho x_{v}\leq 2e(G)\leq 2sn.$
Consequently, $|V_{1}|\leq(5s)^{-6}n$, as desired. Similarly, $|V_{2}|\leq(5s)^{-5}n.$
For any vertex $u\in V(G)$, let $N_{2}(u)$ denote the set of all vertices that are at a distance of two from $u$ in the graph $G$. We can categorize the vertices around $u$ into several sets based on their distances from $u$ and their memberships in $V_{1}$ and $V_{2}$. Specifically, with respect to $V_{1}$, we denote
  • $A_{1}=V_{1}\cap N(u)$, $B_{1}=N(u)\setminus A_{1}$, $A_{2}=V_{1}\cap N_{2}(u)$, $B_{2}=N_{2}(u)\setminus A_{2}$.
Similarly, concerning the set $V_{2}$, we have
  • $C_{1}=V_{2}\cap N(u)$, $D_{1}=N(u)\setminus C_{1}$, $C_{2}=V_{2}\cap N_{2}(u)$, $D_{2}=N_{2}(u)\setminus C_{2}$.
Claim 2.2.
For every $u\in V(G)$, we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}\leq|N(u)|x_{u}+\Big{(}\frac{2sn}{(5s)^{6}}+\frac{2sn}{(5s)^{4}}\Big{)}x_{u^{*}}+\sum_{v\in B_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{w}.$ (1)
Proof.
Given that $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, we have the following inequality for any vertex $u\in V(G)$:
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}\leq\rho^{2}x_{u}=\sum_{v\in N(u)}\rho x_{v}=|N(u)|x_{u}+\sum_{v\in N(u),w\in N(v)\setminus\{u\}}x_{w}.$ (2)
Note that $N(u)=A_{1}\cup B_{1}$. Then $N(v)\setminus\{u\}=N_{A_{1}\cup A_{2}}(v)\cup N_{B_{1}\cup B_{2}}(v)$ for each $v\in N(u)$. By the definition of $V_{1}$, it follows that $x_{w}<(5s)^{-4}x_{u^{*}}$ for every $w\in B_{1}\cup B_{2}$.
Therefore, we can establish an upper bound for the vertices in $A_{1}$ as follows:
$\sum_{v\in A_{1},w\in N(v)\setminus\{u\}}x_{w}$ $=\sum_{v\in A_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{u^{*}}+\sum_{v\in A_{1},w\in N_{B_{1}\cup B_{2}}(v)}(5s)^{-4}x_{u^{*}}$
$\leq\big{(}2e(A_{1})+e(A_{1},A_{2})\big{)}x_{u^{*}}+e(A_{1},B_{1}\cup B_{2})(5s)^{-4}x_{u^{*}}.$ (3)
Similarly, for vertices in $B_{1}$, we obtain
$\sum_{v\in B_{1},w\in N(v)\setminus\{u\}}x_{w}\leq\sum_{v\in B_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{w}+\big{(}2e(B_{1})+e(B_{1},B_{2})\big{)}(5s)^{-4}x_{u^{*}}.$ (4)
Since $G$ satisfies the property $P_{t,r}$ and $s>t+1$, we have that $2e(A_{1})+e(A_{1},A_{2})\leq 2s|V_{1}|$ and $e(A_{1},B_{1}\cup B_{2})+2e(B_{1})+e(B_{1},B_{2})\leq 2sn.$ Additionally, by Claim 2.1, we know that $|V_{1}|\leq(5s)^{-6}n$. Combining these three inequalities with our previous results (2-4), we can derive the desired inequality (1) directly. ∎
Using an argument analogous to that in Claim 2.2, with respect to $V_{2}$, we can find that for every $u\in V(G)$,
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}\leq|N(u)|x_{u}+\frac{4sn}{(5s)^{5}}x_{u^{*}}+\sum_{v\in D_{1},w\in N_{C_{1}\cup C_{2}}(v)}x_{w}.$ (5)
Claim 2.3.
We have $|V_{1}|<(5s)^{6}$.
Proof.
We begin by proving that $|N(u)|\geq(5s)^{-5}n$ for each $u\in V_{1}$. Suppose, for the sake of contradiction, that there exists a vertex $u_{0}\in V_{1}$ with $|N(u_{0})|<(5s)^{-5}n$.
By inequality (5), we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u_{0}}\leq|N(u_{0})|x_{u_{0}}+\frac{4sn}{(5s)^{5}}x_{u^{*}}+\sum_{v\in D_{1},w\in N_{C_{1}\cup C_{2}}(v)}x_{w}.$ (6)
Recall that $D_{1}\subseteq N(u_{0})$ and $C_{1}\cup C_{2}\subseteq V_{2}$. By the property $P_{t,r}$, we know that
$e(D_{1},C_{1}\cup C_{2})\leq e(N(u_{0})\cup V_{2})<s(|N(u_{0})|+|V_{2}|).$
Given our assumption, $|N(u_{0})|<(5s)^{-5}n$, and by Claim 2.1, $|V_{2}|\leq(5s)^{-5}n$. Hence,
$|N(u_{0})|x_{u_{0}}+\sum_{v\in D_{1},w\in N_{C_{1}\cup C_{2}}(v)}x_{w}\leq\Big{(}|N(u_{0})|+e(D_{1},C_{1}\cup C_{2})\Big{)}x_{u^{*}}<\frac{(2s+1)n}{(5s)^{5}}x_{u^{*}}.$
Combining this with inequality (6), we get that $\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u_{0}}<\frac{6s+1}{(5s)^{5}}nx_{u^{*}}.$ On the other hand, since $\lfloor t\rfloor\geq 1$ and $x_{u_{0}}\geq(5s)^{-4}x_{u^{*}}$ (because $u_{0}\in V_{1}$), we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u_{0}}\geq\frac{7}{10}nx_{u_{0}}>\frac{7s}{(5s)^{5}}nx_{u^{*}},$
leading to a contradiction. Therefore, $|N(u)|\geq(5s)^{-5}n$ for each vertex $u\in V_{1}$.
Summing this inequality over all vertices $u\in V_{1}$, we obtain
$|V_{1}|(5s)^{-5}n\leq\sum_{u\in V_{1}}|N(u)|\leq\sum_{u\in V(G)}|N(u)|=2e(G)<2sn.$
Thus, we conclude that $|V_{1}|<(5s)^{6}$. ∎
Claim 2.4.
For each $u\in V_{1}$, we have $|N(u)|\geq\big{(}\frac{x_{u}}{x_{u^{*}}}-\frac{1}{(5s)^{3}}\big{)}n$.
Proof.
Consider $S$ as a subset of vertices from $B_{1}$, with each vertex in $S$ having at least $\lfloor t\rfloor$ neighbors within $A_{1}\cup A_{2}$. Let $p=|A_{1}|+|A_{2}|$. We aim to prove that the size of $S$ is bounded by $|S|\leq s\cdot\ell$, where $\ell=\binom{p}{\lfloor t\rfloor}$.
If $p\leq\lfloor t\rfloor-1$, then it’s impossible for any vertex in $B_{1}$ to have $\lfloor t\rfloor$ neighbors in $A_{1}\cup A_{2}$, which implies that $S$ must be empty, and the statement holds trivially. Therefore, we focus on the case where $p\geq\lfloor t\rfloor$.
Suppose, for contradiction, that $|S|>s\cdot\ell$. Each vertex in $S$ connects to at least $\lfloor t\rfloor$ vertices in $A_{1}\cup A_{2}$. Considering all possible subsets of $A_{1}\cup A_{2}$ of size $\lfloor t\rfloor$, there are $\ell=\binom{p}{\lfloor t\rfloor}$ such subsets. Since each vertex in $S$ has at least $\lfloor t\rfloor$ neighbors, each of these vertices can potentially connect to any of the $\ell$ subsets.
Given that $|S|>s\cdot\ell$, this would mean that on average, each $\lfloor t\rfloor$-vertex subset of $A_{1}\cup A_{2}$ must connect to more than $s$ vertices in $S$. By the pigeonhole principle, there must exist at least one specific $\lfloor t\rfloor$-vertex subset of $A_{1}\cup A_{2}$ that connects to more than $s$ vertices in $S$, implying that this subset has over $s$ common neighbors in $S$. Note that $u\notin A_{1}\cup A_{2}$ and $S\subseteq B_{1}\subseteq N(u)$, then $G$ contains a bipartite subgraph isomorphic to $K_{s,\lfloor t\rfloor+1}$. Since $G$ satisfies the property $P_{t,r}$, we have the inequality
$e(K_{s,\lfloor t\rfloor+1})=s(\lfloor t\rfloor+1)\leq t(s+\lfloor t\rfloor+1)+r.$
Solving for $s$ in the inequality above yields that $s\leq\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t},$ which contradicts the definition of $s$. Therefore, it must be that $|S|\leq s\cdot\ell$.
Given that $p\leq|V_{1}|<(5s)^{6}$, it follows that $\ell$ and $|S|$ are constants. For sufficiently large $n$, we can establish the following bounds.
First, consider the number of edges between $S$ and $A_{1}\cup A_{2}$. We have
$e(S,A_{1}\cup A_{2})\leq|S|(|A_{1}|+|A_{2}|)\leq|S||V_{1}|\leq(5s)^{-6}n.$
On the other hand, by the definition of $S$, any vertex in $B_{1}\setminus S$ has fewer than $\lfloor t\rfloor$ neighbors in $A_{1}\cup A_{2}$. Therefore,
$e(B_{1}\setminus S,A_{1}\cup A_{2})\leq|B_{1}\setminus S|(\lfloor t\rfloor-1)\leq|N(u)|(\lfloor t\rfloor-1).$
Combining these two inequalities, we obtain that
$e(B_{1},A_{1}\cup A_{2})\leq|N(u)|(\lfloor t\rfloor-1)+(5s)^{-6}n.$
It follows that
$\sum_{v\in B_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{w}\leq e\big{(}B_{1},A_{1}\cup A_{2}\big{)}x_{u^{*}}\leq\Big{(}|N(u)|(\lfloor t\rfloor-1)+(5s)^{-6}n\Big{)}x_{u^{*}}.$ (7)
Note that $x_{u}\leq x_{u^{*}}$. Substituting inequality (7) into (1) gives that
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}$ $\leq\Big{(}|N(u)|+\frac{(2s+1)n}{(5s)^{6}}+\frac{2sn}{(5s)^{4}}+|N(u)|(\lfloor t\rfloor-1)\Big{)}x_{u^{*}}$
$\leq\lfloor t\rfloor\Big{(}|N(u)|+\frac{3sn}{(5s)^{4}}\Big{)}x_{u^{*}}.$
Rearranging terms yields that $|N(u)|\geq(n-\lfloor t\rfloor)\frac{x_{u}}{x_{u^{*}}}-\frac{3sn}{(5s)^{4}}.$ For sufficiently large $n$, since $\lfloor t\rfloor\frac{x_{u}}{x_{u^{*}}}\leq\lfloor t\rfloor<\frac{2sn}{(5s)^{4}}$, we conclude that $|N(u)|\geq(\frac{x_{u}}{x_{u^{*}}}-\frac{1}{(5s)^{3}})n,$ as desired. ∎
In the following, we give the proof of Theorem 2.1. Recall that $V=\{v\in V(G):\ x_{v}\geq(5s)^{-1}x_{u^{*}}\}$. By the definition of $V$, it is clear that $u^{*}\in V\subseteq V_{1}$. We will show that $V$ is the desired set in Theorem 2.1.
Proof of Theorem 2.1: We start by proving lower bounds on $x_{u}$ and $|N(u)|$ for each $u\in V$. Suppose, for the sake of contradiction, that there exists $u_{0}\in V$ such that $x_{u_{0}}<(1-\frac{1}{50s^{2}})x_{u^{*}}.$
Since $u_{0}\in V$, we have $x_{u_{0}}\geq\frac{x_{u^{*}}}{5s}.$ By Claim 2.4, we know that
$|N(u^{*})|\geq\Big{(}1-\frac{1}{(5s)^{3}}\Big{)}n\quad\text{and}\quad|N(u_{0})|\geq\Big{(}\frac{1}{5s}-\frac{1}{(5s)^{3}}\Big{)}n.$
Define the following four sets: $A_{1}^{*}=V_{1}\cap N(u^{*})$, $B_{1}^{*}=N(u^{*})\setminus A_{1}^{*}$, $A_{2}^{*}=V_{1}\cap N_{2}(u^{*})$ and $B_{2}^{*}=N_{2}(u^{*})\setminus A_{2}^{*}$. From Claim 2.3, we have $|V_{1}|<(5s)^{6}<\frac{n}{(5s)^{3}}$ for sufficiently large $n$. Thus,
$|B_{1}^{*}|\geq|N(u^{*})|-|V_{1}|\geq\Big{(}1-\frac{2}{(5s)^{3}}\Big{)}n.$
Therefore,
$\big{|}B_{1}^{*}\cap N(u_{0})\big{|}\geq\big{|}B_{1}^{*}\big{|}+\big{|}N(u_{0})\big{|}-n\geq\Big{(}\frac{1}{5s}-\frac{3}{(5s)^{3}}\Big{)}n\geq\frac{22n}{125s}.$ (8)
Given (8), $u_{0}$ must have neighbors in $B_{1}^{*}$. Since $B_{1}^{*}\subseteq N(u^{*})$, $u_{0}$ is within distance two from $u^{*}$, implying $u_{0}\in N(u^{*})\cup N_{2}(u^{*})$. Recalling that $u_{0}\in V\subseteq V_{1}$, we conclude $u_{0}\in A_{1}^{*}\cup A_{2}^{*}$. Using inequality (1) with $u=u^{*}$, we get
$\lfloor t\rfloor(n\!-\!\lfloor t\rfloor)x_{u^{*}}$ $\!\leq\!\Big{(}|N(u^{*})|\!+\!\frac{2sn}{(5s)^{6}}\!+\!\frac{2sn}{(5s)^{4}}\!+\!e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\setminus\{u_{0}\}\big{)}\Big{)}x_{u^{*}}\!+\!e(B_{1}^{*},\{u_{0}\})x_{u_{0}}$
$\!\leq\!\Big{(}|N(u^{*})|\!+\!\frac{2sn}{(5s)^{6}}\!+\!\frac{2sn}{(5s)^{4}}\!+\!e\big{(}B_{1}^{*},A_{1}^{*}\cup A_{2}^{*}\big{)}\Big{)}x_{u^{*}}\!+\!e(B_{1}^{*},\{u_{0}\})(x_{u_{0}}\!-\!x_{u^{*}}),$
where $x_{u_{0}}-x_{u^{*}}<-\frac{x_{u^{*}}}{2(5s)^{2}}$ by our assumption.
From (7), we know that $e(B_{1}^{*},A_{1}^{*}\cup A_{2}^{*})\leq(\lfloor t\rfloor-1)|N(u^{*})|+\frac{n}{(5s)^{6}}.$ Furthermore, for large enough $n$, it easy to see $\lfloor t\rfloor^{2}\leq\frac{n}{(5s)^{6}}.$ This leads to
$\lfloor t\rfloor n$ $\leq\lfloor t\rfloor|N(u^{*})|+\frac{2sn}{(5s)^{4}}+\frac{(2s+2)n}{(5s)^{6}}-\frac{e(B_{1}^{*},\{u_{0}\})}{2(5s)^{2}}$
$<\lfloor t\rfloor n+\frac{2.2sn}{(5s)^{4}}-\frac{e(B_{1}^{*},\{u_{0}\})}{2(5s)^{2}},$
which implies that $e(B_{1}^{*},\{u_{0}\})<\frac{22n}{125s},$ contradicting (8). Therefore, $x_{u}\geq(1-\frac{1}{50s^{2}})x_{u^{*}}$ for each $u\in V$. By Claim 2.4, this also means
$|N(u)|\geq\Big{(}1-\frac{1}{2(5s)^{2}}-\frac{1}{(5s)^{3}}\Big{)}n\geq\big{(}1-\frac{1}{(5s)^{2}}\big{)}n.$
To show $|V|=\lfloor t\rfloor$, assume first that $|V|\geq\lfloor t\rfloor+1$. By the previous inequality, every vertex $u\in V$ has at most $\frac{n}{(5s)^{2}}$ non-neighbors. Correspondingly, any $\lfloor t\rfloor+1$ vertices in $V$ have at least $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}$ common neighbors. Notice that $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}\geq s$. Then $G$ contains a bipartite subgraph isomorphic to $K_{\lfloor t\rfloor+1,s}$. By a similar argument in the proof of Claim 2.4, there is a contradiction since $K_{\lfloor t\rfloor+1,s}$ does not satisfy the property $P_{t,r}$. Therefore, we conclude $|V|\leq\lfloor t\rfloor.$
Next, suppose for contradiction that $|V|\leq\lfloor t\rfloor-1.$ Since $u^{*}\in V\setminus(A_{1}^{*}\cup A_{2}^{*})$, we have $|V\cap(A_{1}^{*}\cup A_{2}^{*})|\leq\lfloor t\rfloor-2.$ This implies that
$e\big{(}B_{1}^{*},V\cap(A_{1}^{*}\cup A_{2}^{*})\big{)}\leq(\lfloor t\rfloor-2)n.$
On the other hand, considering the edges between $B_{1}^{*}$ and $(A_{1}^{*}\cup A_{2}^{*})\setminus V$, we have
$e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\setminus V\big{)}\leq e(G)\leq tn+r<sn.$
By the definition of $V$, we know that $x_{w}<\frac{x_{u^{*}}}{5s}$ for each $w\in(A_{1}^{*}\cup A_{2}^{*})\setminus V$. Now, setting $u=u^{*}$ in inequality (1), we obtain
$\lfloor t\rfloor(n\!-\!\lfloor t\rfloor)x_{u^{*}}$ $\leq\Big{(}|N(u^{*})|\!+\!\frac{2.5sn}{(5s)^{4}}\!+\!e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\cap V\big{)}\Big{)}x_{u^{*}}\!+\!e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\setminus V\big{)}\frac{x_{u^{*}}}{5s}$
$\leq\Big{(}n\!+\!\frac{n}{5}\!+\!(\lfloor t\rfloor\!-\!2)n\!+\!\frac{n}{5}\Big{)}x_{u^{*}}$
$=\big{(}\lfloor t\rfloor\!-\!\frac{3}{5}\big{)}nx_{u^{*}}.$
Dividing both sides by $x_{u^{*}}$ and rearranging terms, we find that
$\lfloor t\rfloor(n-\lfloor t\rfloor)\leq\big{(}\lfloor t\rfloor-\frac{3}{5}\big{)}n.$
Simplifying further, we have $\lfloor t\rfloor^{2}\geq\frac{3}{5}n$, a contradiction. Therefore, $|V|=\lfloor t\rfloor$.
Combining both parts, we establish the desired results, $x_{u}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ and $|N(u)|\geq\big{(}1-\frac{1}{25s^{2}}\big{)}n$ for each $u\in V$. Moreover, $|V|=\lfloor t\rfloor$. The proof of Theorem 2.1 is now complete. ∎
3 Proof of Theorem 1.2
Let $G$ be a graph that achieves the maximal spectral radius in $\mathcal{G}$. It is not hard to verify that $K_{\lfloor t\rfloor,n-\lfloor t\rfloor}\in\mathcal{G}$. Therefore, $\rho(G)\geq\rho(K_{\lfloor t\rfloor,n-\lfloor t\rfloor})=\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$.
Let $V\subseteq V(G)$ be a set defined in accordance with Theorem 2.1. We define $U$ as the set of vertices that are complete to $V$, meaning every vertex in $U$ is adjacent to every vertex in $V$. It is clear that $U$ and $V$ are disjoint sets, i.e., $U\cap V=\emptyset$. We then define $W=V(G)\setminus(U\cup V)$.
Claim 3.1.
We have $|U|\geq\frac{n}{2}$.
Proof.
Set $V=\{v_{1},v_{2},\dots,v_{\lfloor t\rfloor}\}$. By the Inclusion-Exclusion Principle, we have
$|U|=\big{|}\cap_{i=1}^{\lfloor t\rfloor}N(v_{i})\big{|}\geq\sum\limits_{i=1}^{\lfloor t\rfloor}\big{|}N(v_{i})\big{|}-(\lfloor t\rfloor-1)\big{|}\cup_{i=1}^{\lfloor t\rfloor}N(v_{i})\big{|}.$
It follows that
$|U|\geq\Big{(}1-\frac{1}{25s^{2}}\Big{)}n\cdot\lfloor t\rfloor-(\lfloor t\rfloor-1)n=\Big{(}1-\frac{\lfloor t\rfloor}{25s^{2}}\Big{)}n\geq\frac{n}{2},$
as claimed. ∎
Claim 3.2.
We have $W=\emptyset$.
Proof.
Choose a vertex $u\in U\cup W$. By Claim 3.1, let $N_{U}(u)=\{u_{1},u_{2},\dots,u_{d}\}$ and denote by $H$ the subgraph in $G$ induced by the set $V\cup N_{U}(u)\cup\{u\}$. Then $e(H)\geq(\lfloor t\rfloor+1)d$. Since $H$ has the property $P_{t,r}$, we have $e(H)\leq t|H|+r=t(\lfloor t\rfloor+1+d)+r$. Combining the two inequalities, we get that $(\lfloor t\rfloor+1-t)d\leq t(\lfloor t\rfloor+1)+r$ and so
$d_{U}(u)=d\leq\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\leq s-1.$
Recall that $W$ is not complete to $V$, meaning each vertex in $W$ is not adjacent to some vertex in $V$. By Theorem 2.1, every vertex in $V$ is not adjacent to at most $\frac{n}{25s^{2}}$ vertices of $G$. Therefore, we have $|W|\leq\frac{\lfloor t\rfloor n}{25s^{2}}$. Note that $d(u)\leq d_{U}(u)+|V|+d_{W}(u)\leq s-1+\lfloor t\rfloor+d_{W}(u)$. Moreover, since $G$ has the property $P_{t,r}$, we have either $e(W)\leq t|W|+r$ or $|W|<t$. Consequently, for $w\in W$, we get that
$\sum\limits_{w\in W}\rho x_{w}\leq\Big{(}\big{(}s-1+\lfloor t\rfloor\big{)}|W|+2e(W)\Big{)}x_{u^{*}}\leq\frac{(s+3t)\lfloor t\rfloor n}{25s^{2}}x_{u^{*}}.$
It follows that
$\sum\limits_{w\in N_{W}(v)}x_{w}\leq\sum\limits_{w\in W}x_{w}<\frac{(s+3t)\lfloor t\rfloor n}{25\rho s^{2}}x_{u^{*}}.$ (9)
Now for $u\in U\cup W$, we have $\rho x_{u}\leq(s-1+\lfloor t\rfloor)x_{u^{*}}+\sum_{w\in N_{W}(u)}x_{w}$. Combining with the inequality (9) and $\rho\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, we deduce that
$x_{u}<\Big{(}\frac{s-1+\lfloor t\rfloor}{\rho}+\frac{(s+3t)\lfloor t\rfloor n}{25\rho^{2}s^{2}}\Big{)}x_{u^{*}}\leq\frac{s+4t}{25s^{2}}x_{u^{*}}<\frac{x_{u^{*}}}{5s}.$ (10)
Now, suppose that $W\neq\emptyset$. Let $G^{*}$ be a graph obtained from $G$ by
  • Removing all edges within $G[W]$.
  • Removing all edges between $U$ and $W$.
  • Adding edges between $V\setminus N(W)$ and $W$.
Additionally, let $w_{1}$ be a vertex of minimal degree in $G[W]$, and let $E_{1}$ be the set of edges incident to $w_{1}$ in $G[W]$. Following the same approach, let $w_{2}$ be a vertex of minimal degree in $G[W\setminus\{w_{1}\}]$, and let $E_{2}$ be the set of edges incident to $w_{2}$ in $G[W\setminus\{w_{1}\}]$. Repeating the process, we define $W_{1}=W$ and $W_{i}=W\setminus\{w_{1},\dots,w_{i-1}\}$ for $i\geq 2$, where $w_{i}$ is a vertex of minimum degree in $G[W_{i}]$ and $E_{i}$ is the set of edges that are incident to $w_{i}$ in $G[W_{i}]$. By continuing this process iteratively, we have
$E(G[W])=\cup_{i=1}^{|W|}E_{i}=\cup_{i=1}^{|W|}\{w_{i}w:~{}w\in N_{W_{i}}(w)\}.$
Recall that $G\in\mathcal{G}$ and $s=\big{\lceil}\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\big{\rceil}+1\geq t^{2}+r+1$. When $|W_{i}|\geq t$, we have
$d_{W_{i}}(w_{i})\leq\frac{2e(W_{i})}{|W_{i}|}\leq\frac{2t|W_{i}|+2r}{|W_{i}|}\leq 2t+\frac{2r}{t}<2s.$
When $|W_{i}|<t$, we also get $d_{W_{i}}(w_{i})\leq\frac{2\binom{|W_{i}|}{2}}{|W_{i}|}<t<2s.$ Now, we can drive that
$\rho(G^{*})-\rho(G)$ $\geq X^{T}\big{(}A(G^{*})-A(G)\big{)}X$
$=\sum\limits_{i=1}^{|W|}2x_{w_{i}}\Big{(}\sum\limits_{v\in V\setminus N_{V}(w_{i})}x_{v}-\sum\limits_{u\in N_{U}(w_{i})}x_{u}-\sum\limits_{w\in N_{W_{i}}(w_{i})}x_{w}\Big{)}.$ (11)
In view of the definition of $W$, we have $|V\setminus N_{V}(w_{i})|\geq 1$ for each $w_{i}\in W$. By Theorem 2.1, $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for each $v\in V$. Moreover, by previous discussion, we know that $d_{U}(u)\leq s-1$ for each $u\in U$ and $d_{W_{i}}(w_{i})<2s$ for each $w_{i}\in W_{i}$. Combining (10) and (3), we can see that
$\rho(G^{*})-\rho(G)\geq\sum\limits_{w_{i}\in W}2x_{w_{i}}\Big{(}\big{(}1-\frac{1}{50s^{2}}\big{)}-\big{(}3s-1\big{)}\frac{1}{5s}\Big{)}x_{u^{*}}>0.$
By the maximality of $G$, the modified graph $G^{*}$ cannot satisfy the property $P_{t,r}$. Specifically, there exists an induced subgraph $H$ of $G^{*}$ such that $|H|\geq t$ and $e(H)>t|H|+r$. Since $G^{*}[U\cup V]=G[U\cup V]$, the violation of the property $P_{t,r}$ must involve vertices in $W$, i.e., $V(H)\cap W\neq\emptyset$.
Let $H^{\prime}=G^{*}[V(H)\setminus W]$. According to the definition of $G^{*}$, the size of $H^{\prime}$ can be expressed as: $e(H)=e(H^{\prime})+|V(H)\cap W|\cdot|V(H)\cap V|>t|V(H)|+r.$ Since $|V(H)\cap V|\leq\lfloor t\rfloor\leq t$, it follows that $e(H^{\prime})>t|V(H)|+r-t|V(H)\cap W|=t|V(H^{\prime})|+r$.
If $|V(H^{\prime})|\geq t$, this inequality leads to a contradiction because $G^{*}[V(H^{\prime})]=G[V(H^{\prime})]$, and $G$ is assumed to satisfy the property $P_{t,r}$.
If $|V(H^{\prime})|<t$, given that $|U\setminus V(H^{\prime})|\geq n/2-t\geq|W|$, there are sufficiently many vertices in $U$ to replace $V(H)\cap W$. We can define $H^{\prime\prime}=G^{*}[U^{\prime}\cup V(H^{\prime})]$, where $U^{\prime}\subseteq U\setminus V(H^{\prime})$ and $|U^{\prime}|=|W|$. Now, the size of $H^{\prime\prime}$ can be expressed as: $e(H^{\prime\prime})=e(H)+e(V(H)\cap U,U^{\prime})+e(U^{\prime})>t|V(H)|+r=t|V(H^{\prime})|+r$. This also leads to a contradiction.
In both cases, we arrive at a contradiction, which implies that our initial assumption $W\neq\emptyset$ must be false. Therefore, $W=\emptyset$. ∎
By Claim 3.2, we have $V(G)=V\cup U$ and $e(V,U)=|V||U|$, where $|V|=\lfloor t\rfloor$. Recall that every $u\in U$ has at most $s-1$ neighbors in $U$. Let $u^{\star}\in U$ such that $x_{u^{\star}}=\max_{u\in U}x_{u}$. Then we have $\rho x_{u^{\star}}\leq(s-1)x_{u^{\star}}+\lfloor t\rfloor x_{u^{*}}.$ It follows that $x_{u^{\star}}\leq\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}.$ Also, since $u^{*}\in V$, we have $\rho x_{u^{*}}\leq(\lfloor t\rfloor-1)x_{u^{*}}+(n-\lfloor t\rfloor)x_{u^{\star}}.$ Combining these two inequalities, we get that
$(\rho-\lfloor t\rfloor+1)x_{u^{*}}\leq(n-\lfloor t\rfloor)\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}.$
Dividing both sides by $x_{u^{*}}$, we obtain $\rho-\lfloor t\rfloor+1\leq\frac{(n-\lfloor t\rfloor)\lfloor t\rfloor}{\rho-s+1}.$ Expanding and rearranging terms, we have
$\rho^{2}-(s+\lfloor t\rfloor-2)\rho-\lfloor t\rfloor n+\lfloor t\rfloor^{2}+(s-1)(\lfloor t\rfloor-1)\leq 0.$
Solving this quadratic inequality, we find that
$\rho\leq\frac{(s+\lfloor t\rfloor-2)}{2}+\sqrt{\frac{(s+\lfloor t\rfloor-2)^{2}}{4}+\lfloor t\rfloor n-\lfloor t\rfloor^{2}-(s-1)(\lfloor t\rfloor-1)}.$
Therefore, $\rho\leq c(s,t)+\sqrt{\lfloor t\rfloor n},$ where $c(s,t)$ is a constant depending on $s$ and $t$. This completes the proof of Theorem 1.2.
4 Extremal graphs
In this section, we will prove Theorem 1.3 and characterize the structure of the extremal graph. Before delving into our proof, we first introduce some concepts and some key lemmas. For a positive integer $\ell$, an $\ell$-walk in a graph $G$ consists of $\ell+1$ vertices where consecutive vertices are adjacent. Let $w_{G}^{\ell}(u)$ be the number of $\ell$-walks starting at vertex $u$ in $G$. The total number of $\ell$-walks in $G$ is defined as $W_{\ell}(G)=\sum_{v\in V(G)}w_{G}^{\ell}(v)$. Obviously, $W_{1}(G)=2e(G)$ and $W_{2}(G)=\sum_{v\in V(G)}(d_{G}(v))^{2}.$
Let $G_{1}$ and $G_{2}$ be two graphs, possibly of different orders. We say $G_{1}$ is strictly walk-preferable to $G_{2}$, if there exists a positive integer $\ell$ such that $W_{\ell}(G_{1})>W_{\ell}(G_{2})$ and $W_{i}(G_{1})=W_{i}(G_{2})$ for every $i\leq\ell-1$. Otherwise, $W_{\ell}(G_{1})=W_{\ell}(G_{2})$ for all $\ell\geq 1$, then we say $G_{1}$ is walk-equivalent to $G_{2}$ and denote $G_{1}\equiv G_{2}$. We write $G_{2}\prec G_{1}$ if $G_{1}$ is strictly walk-preferable to $G_{2}$, and we write $G_{2}\preceq G_{1}$ if $G_{1}$ is walk-preferable to $G_{2}$ (either strictly or equivalently).
The following theorem, provided by Zhang, offers a framework for comparing the spectral radius of graphs based on their local walk structures. This seems particularly useful in various graph-theoretic analyses and applications.
Lemma 4.1.
([21]) Let $G$ be a connected graph of order $n$. Assume that $S$ is a non-empty proper subset of $V(G)$ and $T$ is a set of isolated vertices in $G-S$ such that $e(S,T)=|S||T|$. Let $H_{1},H_{2}$ be two graphs with the same vertex set $T$, and let $G_{i}$ be the graph obtained from $G$ by embedding the edges of $H_{i}$ into $T$. When $\rho(G)$ is sufficiently large compared to $|T|$, we have the following conclusions:
  • (i)
    If $H_{1}\equiv H_{2}$, then $\rho(G_{1})=\rho(G_{2})$.
  • (ii)
    If $H_{1}\succ H_{2}$, then $\rho(G_{1})>\rho(G_{2})$.
  • (iii)
    If $H_{1}\prec H_{2}$, then $\rho(G_{1})<\rho(G_{2})$.
Let $\mathbb{G}(n,m)$ be the family of graphs with $n$ vertices and $m$ edges, and let $F_{p}(H)$ be the sum of the $p$-th powers of vertex degrees of a graph $H$, formally defined as
$F_{p}(H)=\sum_{v\in V(H)}\big{(}d_{H}(v)\big{)}^{p},$
where $d_{H}(v)$ is the degree of vertex $v$ in $H$, and $p\geq 1$ is an integer. Ismailescu and Stefanica showed the following result regarding the maximization of $F_{p}(H)$.
Lemma 4.2.
([14]) Let $H^{*}\in\mathbb{G}(n,m)$ such that $F_{p}(H^{*})=\max_{H\in\mathbb{G}(n,m)}F_{p}(H)$, where $p\geq 2$ is an integer. If $n\geq m+2$, then $H^{*}$ is the star $K_{1,m}$ plus $n-m-1$ isolated vertices. Moreover, it is unique except in the case when $p=2$ and $m=3$ (when both $K_{1,3}\cup(n-4)K_{1}$ and $K_{3}\cup(n-3)K_{1}$ are extremal graphs).
Considering that $p=2$ and $n=m-1$, similarly to Lemma 4.2 one can get the following result. For ease of understanding and to maintain integrity, we present the full proof here.
Lemma 4.3.
Let $H^{*}\in\mathbb{G}(n,m)$ such that $F_{2}(H^{*})=\max_{H\in\mathbb{G}(n,m)}F_{2}(H)$. If $m=n-1$, then $H^{*}\cong K_{1,m}$. Moreover, this configuration is unique unless $m=3$, in which case both $K_{1,3}$ and $K_{3}\cup K_{1}$ are extremal graphs.
Proof.
Consider a graph $H^{*}\in\mathbb{G}(n,m)$ maximizing $F_{2}(H)$ over all graphs with $n$ vertices and $m$ edges. Let $(d_{1},d_{2},...,d_{n})$ represent the degrees of vertices $v_{1},v_{2},...,v_{n}$ in $H^{*}$. We first aim to show that the vertex of maximal degree in $H^{*}$ connects to all non-isolated vertices.
Suppose there exists a vertex $v_{i}$ not connected to another vertex $v_{j}$ with $d_{i}\leq d_{j}$, but both are connected to a third vertex $v_{k}$. By replacing the edge $v_{i}v_{k}$ with $v_{j}v_{k}$, we form a new graph $H^{\prime}\in\mathbb{G}(n,m)$. Then we have
$F_{2}(H^{\prime})-F_{2}(H^{*})=(d_{j}+1)^{2}+(d_{i}^{2}-1)-d_{j}^{2}-d_{i}^{2}=2d_{j}-2d_{i}+2>0,$
contradicting the maximality of $H^{*}$. Thus, the vertex of maximal degree must connect to all other non-isolated vertices.
Assume $v_{n}$ has maximal degree $d_{n}=l$ and is adjacent to $l$ vertices $\{v_{1},v_{2},...,v_{l}\}$. Since $m=n-1$, if $l<n-1$, it implies that $d_{l+1}=\cdots=d_{n-1}=0$. Our goal is to prove that $l=n-1$, which would make $H^{*}$ a star $K_{1,m}$.
Let $1\leq d_{1}\leq d_{2}\leq\cdots\leq d_{l}\leq d_{n}=l$. Suppose for contradiction that $l<n-1$. Then the subgraph induced by $\{v_{1},v_{2},...,v_{l}\}$ contains $n-1-l$ edges, it implies that $d_{l}>1$ and $d_{l}-1\leq n-1-l$.
We construct a new graph $H^{\prime\prime}$ from $H^{*}$ by removing the $d_{l}-1$ edges corresponding to $v_{l}$ from the subgraph with vertices $v_{1},\ldots,v_{l}$ and replacing them with $d_{l}-1$ edges between $v_{n}$ and $d_{l}-1$ isolated vertices. Note that we have sufficiently many isolated vertices to do this construction, since $d_{l}-1\leq n-1-l$.
The vertex degrees of $H^{\prime\prime}$ have changed as follows: $d^{\prime}_{l}=1,d^{\prime}_{l+1}=\cdots=d^{\prime}_{l+d_{1}-1}=1,d^{\prime}_{n}=l+d_{1}-1$, and for $11\leq i\leq l-$ 1, $d^{\prime}_{i}=d_{i}-1$ for $d_{l}-1$ vertices and $d^{\prime}_{i}=d_{i}$ for the others. We recall that $d_{i}\leq l$ for $1\leq i\leq l$. By direct computation, we obtain
$F_{2}(H^{*})-F_{2}(H^{\prime\prime})=\sum_{i=1}^{l-1}(d_{i}^{2}-d_{i}^{\prime 2})-(d_{l}-1)(2l-1).$
If $d_{i}^{\prime}=d_{i}-1$, then $d_{i}^{2}-d_{i}^{\prime 2}=2d_{i}-1\leq 2l-1$. It results that
$\sum_{i=1}^{l-1}(d_{i}^{2}-d_{i}^{\prime 2})\leq(d_{l}-1)(2l-1).$
Thus, we obtain that $F_{2}(H^{*})-F_{2}(H^{\prime\prime})\leq 0$ and $F_{2}(H^{*})-F_{2}(H^{\prime\prime})=0$ if and only if, for all indices $11\leq i\leq l-$ 1 for which $d_{i}^{\prime}=d_{i}-1$, we have $d_{i}=l$. Recall that $d_{i}\leq d_{l}$. Thus $d_{l}=l$, and this implies that all the vertices $v_{1},\ldots,v_{l-1}$, are neighbors of $v_{l}$. By the above observation, it results that $d_{1}=d_{2}=\cdots=d_{l}=l.$
Therefore, the only scenario in which our delete-add construction (performed under the assumption $d_{l}>1$) would not result in a graph $H^{\prime\prime}$ with a strictly higher $F_{2}$ value than our initial graph $H^{*}$ is when the only nontrivial component of $H^{*}$ is a complete graph on $l+1$ vertices.
In this case, we can compare directly $H=K_{l+1}\cup(n-1-l)K_{1}$ with our candidate $H^{*}=K_{1,m}$. We note that $m=\binom{l+1}{2}$, and obtain
$F_{2}(H)=(l+1)l^{2}\quad\text{and}\quad F_{2}(H^{*})=\binom{l+1}{2}^{2}+\binom{l+1}{2}.$
This implies that
$F_{2}(H^{*})-F_{2}(H)=\frac{(l-2)(l-1)l(l+1)}{4}\geq 0.$
This difference is non-negative, with equality holding only when $l=1$ or $l=2$. For $l=1$, $H^{*}$ becomes $K_{1,1}=K_{2}$, and for $l=2$, $H^{*}$ can also be $K_{3}\cup K_{1}$ when $m=3$.
Therefore, under the conditions specified, $H^{*}$ is uniquely the star $K_{1,m}$, except for the special case when $m=3$ where $K_{3}\cup K_{1}$ is also an extremal graph. ∎
We now formally provide the proof of Theorem 1.3. Let $G$ be a spectral extremal graph that satisfies the property $P_{t,r}$. By Claim 3.2, we know that $V(G)=V\cup U$, where $|V|=\lfloor t\rfloor$ and $e(V,U)=|V||U|$. Let $T$ be a subset of $U$ such that $e(T,U\setminus T)=0$ and $|T|$ is constant. Consider two graphs $H_{1}$ and $H_{2}$ with $V(H_{1})=V(H_{2})=T$ and $e(H_{1})=e(H_{2})$. Define $G_{i}$ as the graph obtained from $G$ by embedding the edges of $H_{i}$ into $T$ for $i\in\{1,2\}$. Based on Lemma 4.1, we have the following observation.
Fact 4.1.
If $H_{1}\succ H_{2}$ (specially, $F_{2}(H_{1})>F_{2}(H_{2})$), then $\rho(G_{1})>\rho(G_{2})$.
We begin by establishing the following key claim.
Claim 4.1.
$V$ is a clique.
Proof.
Suppose, for the sake of contradiction, that there exist two vertices $v_{1},v_{2}\in V$ such that $v_{1}v_{2}\not\in E(G)$. Since $\rho(G+v_{1}v_{2})>\rho(G)$, $G$ contains a subgraph $H$ such that $H+v_{1}v_{2}$ no longer satisfies the property $P_{t,r}$. We call $H+v_{1}v_{2}$ a counterexample of $G+v_{1}v_{2}$. Without loss of generality, we may assume that $H+v_{1}v_{2}$ is a minimal counterexample in terms of the number of vertices.
Let $A=V\cap V(H)$, $B=U\cap V(H)$ and $C=V\setminus V(H)$. Define $L$ as the vertex set of all components in $G[U\setminus B]$ that have at least one edge connecting to some vertex in $B$. Furthermore, let $N_{0}$ denote the vertex set of all tree-components in $G[U\setminus(B\cup L)]$ and $N_{1}$ be the vertex set of all other components in $G[U\setminus(B\cup L)]$.
First, we will show that $L\cup N_{1}$ is finite. Consider the size of the induced subgraph $G[V\cup B\cup L\cup N_{1}]$:
$e(V\cup B\cup L\cup N_{1})=e(H)+e(C,V(H))+e(C)+e(V,L\cup N_{1})+e(B,L)+e(L)+e(N_{1}).$
By the property $P_{t,r}$, we have
$e(V\cup B\cup L\cup N_{1})\leq t(|V|+|L|+|N_{1}|)+r.$
Note that $e(V(H))=\lfloor t|V(H)|+r\rfloor$, $e(B,L)+e(L)\geq|L|$, $e(N_{1})\geq|N_{1}|$. Therefore, it follows that
$e(V\cup B\cup L\cup N_{1})\geq\lfloor t|V(H)|+r\rfloor+|L|+|N_{1}|+(|L|+|N_{1}|)\lfloor t\rfloor.$
Combining the above two inequalities, we obtain
$(\lfloor t\rfloor+1-t)(|L|+|N_{1}|)$ $\leq t|V(H)|+r-\lfloor t|V(H)|+r\rfloor+t|C|$
$\leq 1+t|C|\leq 1+t\lfloor t\rfloor,$
which shows that $|L|+|N_{1}|$ is finite.
Now, we partition $N_{0}$ into two parts: Let $N_{0}^{1}$ denote the vertex set of all components with cardinality at least $1/t_{0}$, let $N_{0}^{2}$ denote the vertex set of all remaining components, where $t_{0}=\lfloor t\rfloor+1-t$.
Using a similar argument as before, we can show that $N_{0}^{1}$ is also finite. Consider the graph $G[V\cup B\cup N_{0}^{1}]$, we have
$e(V\cup B\cup N_{0}^{1})=e(V(H))+e(C,V(H))+e(C)+e(N_{0}^{1},V)+e(N_{0}^{1}).$
Given that $e(N_{0}^{1})$ for a forest $G[N_{0}^{1}]$ is at most $|N_{0}^{1}|-1$, and using the property $P_{t,r}$, we have
$e(V\cup B\cup N_{0}^{1})$ $\leq\lfloor t|V(H)|+r\rfloor+e(C,V(H))+e(C)+\lfloor t\rfloor|N_{0}^{1}|+|N_{0}^{1}|-1$
$\leq t(|V(H)|+|C|+|N_{0}^{1}|)+r.$
Since $20\leq|C|\leq\lfloor t\rfloor-$ 2, it follows that
$t_{0}|N_{0}^{1}|\leq t|C|-e(C,V(H))-e(C)+2<t^{2}.$
If $C\neq\emptyset$, we consider the graph $G[V\cup B]$:
$e(V\cup B)=e(V(H))+e(C,A)+e(C)+e(C,B).$
Using the property $P_{t,r}$, we get
$e(V\cup B)$ $=\lfloor t|V(H)|+r\rfloor+e(C,A)+e(C)+|C||B|$
$\leq t(|V(H)|+|C|)+r.$
Since $21\leq|C|\leq\lfloor t\rfloor-$ 2, we obtain
$|C||B|\leq t|C|-e(C,A)-e(C)+1\leq t|C|+1,$
which implies that
$|B|\leq t+1.$
Let $G^{\prime}$ be formed from $G$ by removing all edges induced by $B\cup L\cup N_{1}\cup N_{0}^{1}$ in $G$ and adding the edge $v_{1}v_{2}$. Since each component in $N_{0}^{2}$ has cardinality less than $1/t_{0}$, it is straightforward to verify that $G^{\prime}$ satisfies the property $P_{t,r}$.
According to the proof of Theorem 2.1, we have $x_{u}\leq\frac{\lfloor t\rfloor}{\rho-s}x_{u^{*}}$ for $u\in U$, and $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for $v\in V$. Let $|B\cup L\cup N_{1}\cup N_{0}^{1}|=c$. Note that $\rho\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$. Then, we can deduce that
$\rho(G^{\prime})-\rho(G)$ $\geq X^{T}(A(G^{\prime})-A(G))X$
$=2x_{v_{1}}x_{v_{2}}-2\sum_{u_{i},u_{j}\in U\setminus N_{0}^{2}}x_{u_{i}}x_{u_{j}}$
$\geq 2\big{(}1-\frac{1}{50s^{2}}\big{)}^{2}x_{u^{*}}^{2}-2\big{(}\frac{c\lfloor t\rfloor}{\rho-s}\big{)}^{2}x_{u^{*}}^{2}$
$=2\big{[}\big{(}1-\frac{1}{50s^{2}}\big{)}^{2}-\big{(}\frac{c\lfloor t\rfloor}{\rho-s}\big{)}^{2}\big{]}x_{u^{*}}^{2}$
$>0,$
which is a contradiction since $G$ was assumed to have the maximum spectral radius.
If $C=\emptyset$, then we construct a new graph $G^{\prime\prime}$ as follows:
  • Delete all edges induced by $L\cup N_{1}\cup N_{0}^{1}$.
  • Delete all edges between $B$ and $L$.
  • Delete one edge in $G[B]$.
  • Add the edge $v_{1}v_{2}$.
The construction of $G^{\prime\prime}$ ensures that it still satisfies the property $P_{t,r}$ while potentially increasing the spectral radius, leading to a similar contradiction. Consequently, the contradictions derived from both cases force us to conclude that $V$ must form a clique, as desired. ∎
With Claim 4.1 established, we can classify the components of $G[U]$ into three distinct categories:
  • $\mathcal{N}_{1}$ includes all components that are not trees.
  • $\mathcal{N}_{2}$ contains all components that are nontrivial trees.
  • $\mathcal{N}_{3}$ consists solely of isolated vertices.
Moreover, we denote
$g_{0}=t|V|+r-e(V),$
and for a graph $H$ with $V(H)\subseteq U$, we define
$\eta(H)=e(H)+(\lfloor t\rfloor-t)|V(H)|.$
Let $F\subseteq G[U]$ and $H$ is a component of $F$, we denote that
$\eta^{+}(F)=\sum_{\eta(H)>0}\eta(H).$
Obviously, $g_{0}$ is nonnegative and constant as $r\geq-\binom{\lfloor t+1\rfloor}{2}$ and $V$ is a $\lfloor t\rfloor$-clique. By the definitions of $g_{0}$ and $\eta(H)$, we immediately obtain the following observation.
Fact 4.2.
If $K_{|V|}\nabla H$ satisfies the property $P_{t,r}$, then $\eta(H)\leq g_{0}$. Moreover, if $F$ is a forest with $\eta^{+}(F)\leq g_{0}$, then $K_{|V|}\nabla F$ satisfies the property $P_{t,r}$.
Given that each component $H$ in $\mathcal{N}_{1}$ has at least $|V(H)|$ edges, we have $e(H)\geq|V(H)|$. Thus, for any such component $H$:
$\eta(H)=e(H)+\lfloor t\rfloor|V(H)|-t|V(H)|\geq|V(H)|+\lfloor t\rfloor|V(H)|-t|V(H)|\geq t_{0}|V(H)|.$
Since $t_{0}>0$, it follows that $\eta(H)>0$. Summing over all components $H\in\mathcal{N}_{1}$, we have
$t_{0}\sum_{H\in\mathcal{N}_{1}}|V(H)|\leq\sum_{H\in\mathcal{N}_{1}}\eta(H)\leq g_{0}.$
Therefore,
$\sum_{H\in\mathcal{N}_{1}}|V(H)|\leq\frac{g_{0}}{t_{0}}.$
This inequality indicates that the total number of vertices across all components in $\mathcal{N}_{1}$ is bounded by $\frac{g_{0}}{t_{0}}$.
For a component $H$ in $\mathcal{N}_{2}$, we have $e(H)=|V(H)|-1$. Thus, the function $\eta(H)$ for such a component becomes
$\eta(H)=(|V(H)|-1)+\lfloor t\rfloor|V(H)|-t|V(H)|=t_{0}|V(H)|-1.$
For $\eta(H)>0$,
$t_{0}|V(H)|-1>0\implies|V(H)|>\frac{1}{t_{0}}.$
This implies that components with $\eta(H)>0$ must have more than $\frac{1}{t_{0}}$ vertices. And similar to above argument, the total number of vertices of these components is bounded by $\frac{g_{0}+1}{t_{0}}$. Conversely, if $\eta(H)\leq 0$, then
$t_{0}|V(H)|-1\leq 0\implies|V(H)|\leq\frac{1}{t_{0}}.$
Such components have a limited number of vertices, which is at most $\lfloor\frac{1}{t_{0}}\rfloor$. Therefore, the remaining components, having $\eta(H)\leq 0$, contain lots of vertices that is $\Theta(n)$.
Claim 4.2.
Each component $H$ in $\mathcal{N}_{2}$ is a star.
Proof.
Let $H\in\mathcal{N}_{2}$ be a tree. Then $e(H)=|V(H)|-1$. According to Lemma 4.3, the star is walk-preferable to other trees with the same size. And by Lemma 4.1, a graph which has a walk-preferable subgraph in a same vertex set has a greater spectral radius. Since $G$ has a maximal spectral radius, $H$ must be a star. ∎
Having characterized the components in $\mathcal{N}_{2}$ as stars, we next examine the family $\mathcal{N}_{1}$.
Claim 4.3.
The family $\mathcal{N}_{1}$ contains at most one component.
Proof.
We aim to show that the family $\mathcal{N}_{1}$ contains at most one component. To do this, we will proceed by contradiction. Suppose there are two distinct components $H_{1}$ and $H_{2}$ in $\mathcal{N}_{1}$. Since each component in $\mathcal{N}_{1}$ contains at least one cycle, both $H_{1}$ and $H_{2}$ must have at least as many edges as vertices. Therefore, we get
$e(H_{1})+e(H_{2})\geq|V(H_{1})|+|V(H_{2})|\geq 6.$
Define $a=e(H_{1})+e(H_{2})-|V(H_{1})|-|V(H_{2})|+1$. Now, we consider the graph $H$ formed by the union of $H_{1}$, $H_{2}$, and $a$ copies of the star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ (note that there are sufficient many such stars.):
$H=H_{1}\cup H_{2}\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}.$
If $\lfloor\frac{1}{t_{0}}\rfloor=1$, then $H$ simplifies to
$H=H_{1}\cup H_{2}\cup aK_{1}.$
In either case, it is clear that
$e(H)=|V(H)|-1.$
Transform $H$ into a star $H^{\prime}=K_{1,|V(H)|-1}$. Starting from the original graph $G$, construct a new graph $G^{\prime}$ by removing the edges of $H$ and adding the edges of $H^{\prime}$. Note that $e(H)=e(H^{\prime})$ and $V(H)=V(H^{\prime})$. The parameter $\eta$ for $H^{\prime}$ is calculated as follows:
$\eta(H^{\prime})=e(H^{\prime})-(1-t_{0})|V(H^{\prime})|.$
Substituting the values:
$\eta(H^{\prime})$ $=e(H_{1})+e(H_{2})+a(\lfloor\frac{1}{t_{0}}\rfloor-1)-(1-t_{0})(|V(H_{1})|+|V(H_{2})|+a\lfloor\frac{1}{t_{0}}\rfloor)$
$=e(H_{1})-(1-t_{0})|V(H_{1})|+e(H_{2})-(1-t_{0})|V(H_{2})|-a+at_{0}\lfloor\frac{1}{t_{0}}\rfloor$
$=\eta(H_{1})+\eta(H_{2})-a(1-t_{0}\lfloor\frac{1}{t_{0}}\rfloor).$
Since $1-t_{0}\lfloor\frac{1}{t_{0}}\rfloor\geq 0$ and $a\geq 0$, it follows that
$\eta(H^{\prime})\leq\eta(H_{1})+\eta(H_{2})\leq g_{0}.$
By Fact 4.2, $G^{\prime}$ satisfies the property $P_{t,r}$. By Lemma 4.3, we have $F_{2}(H)<F_{2}(H^{\prime})$. According to Fact 4.1, we get $\rho(G)<\rho(G^{\prime})$. This contradicts the assumption that $G$ was optimal (or maximal). Therefore, $\mathcal{N}_{1}$ can contain at most one component. ∎
Having constrained the number of non-tree components in $\mathcal{N}_{1}$, we now turn our attention to the tree components in $\mathcal{N}_{2}$, particularly those that are stars. We aim to show that there can be at most one star of a specific size in $\mathcal{N}_{2}$.
Claim 4.4.
$\mathcal{N}_{2}$ has at most one star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$ with $a\geq 1$.
Proof.
Suppose, for contradiction, that there are two stars $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}}$ and $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}}$ in $\mathcal{N}_{2}$ with $n_{1}\geq n_{2}\geq 1$. Construct a new graph $G^{\prime}$ from $G$ by a edge-shifting operation, i.e., replacing $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}}$ with $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1}$. Note that
$\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1})+\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1})=\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}})+\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}})\leq g_{0},$
and
$F_{2}(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}})<F_{2}(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1}).$
Since $(\lfloor\frac{1}{t_{0}}\rfloor+1)t_{0}-1>\frac{1}{t_{0}}t_{0}-1=0$ and $n_{2}\geq 1$, we have
$g_{0}\geq\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1})=(\lfloor\frac{1}{t_{0}}\rfloor+n_{2})t_{0}-1=(\lfloor\frac{1}{t_{0}}\rfloor+1)t_{0}-1+(n_{2}-1)t_{0}>0.$
It follows that
$\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1})\leq g_{0}-\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1})<g_{0}.$
Consequently, $G^{\prime}$ satisfies the property $P_{t,r}$ and $\rho(G^{\prime})>\rho(G)$, a contradiction. This completes the proof. ∎
With the constraints on $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ established, we can now link these findings to the overall structure of $G[U]$. Specifically, we investigate the implications of having a single non-tree component in $\mathcal{N}_{1}$.
Claim 4.5.
The following statements are equivalent:
  1. (1)
    $|\mathcal{N}_{1}|=1$.
  2. (2)
    $43t_{0}\leq g_{0}<8t_{0}-$ 4.
  3. (3)
    $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
Proof.
$(1)\Rightarrow(2)$: Assume $\mathcal{N}_{1}=\{H_{1}\}$, meaning there is exactly one non-tree component in $G[U]$, which implies $|V(H_{1})|\geq 3$. We will show that $H_{1}\cong K_{3}$.
If $e(H_{1})\geq 4$, we proceed similarly to the proof of Claim 4.3. Consider replacing $H:=H_{1}\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ with a star $H^{\prime}:=K_{1,|V(H)|-1}$, where $a=e(H_{1})+1-|V(H_{1})|$. We can show that this replacement satisfies the property $P_{t,r}$. By Lemmas 4.1 and 4.3, this replacement would yield a new graph with a greater spectral radius than $G$, leading to a contradiction. Therefore, $e(H_{1})=3$.
Given $e(H_{1})=3$ and $|V(H_{1})|\geq 3$, $H_{1}$ must be either $K_{3}$ or $K_{1,3}$ (a claw). Since $W_{3}(K_{3})=24>W_{3}(K_{1,3})=18$, $K_{3}$ is walk-preferable to the star $K_{1,3}$. Thus, $H_{1}\cong K_{3}$.
It follows that $\eta(H_{1})\geq 3t_{0}$, confirming that $H_{1}$ is indeed a triangle.
Next, consider $\eta(4K_{2})=8t_{0}-4$. If $g_{0}\geq 8t_{0}-4$, then $e(U)\geq 4$, as $4K_{2}$ is walk-preferable to $K_{3}\cup 5K_{1}$. This would imply the existence of a star $K_{1,a}$ in $\mathcal{N}_{2}$ with $a\geq 1$. Replacing $K_{3}\cup K_{1,a}$ with $K_{1,3+a}$ would again result in a graph with a greater spectral radius, contradicting our assumption. Therefore, $g_{0}<8t_{0}-4$.
$(2)\Rightarrow(3)$: Given $43t_{0}\leq g_{0}<8t_{0}-$ 4, the upper bound on $g_{0}$ restricts the number of edges in $U$. Specifically, since $\eta(H)\geq 8t_{0}-4$ for any graph $H$ with $e(H)=4$, it follows that $e(U)\leq 3$. Moreover, because $\eta(K_{3})=3t_{0}\leq g_{0}$, we conclude that $e(U)=3$. By Lemma 4.1 and previous discussions, the only graph that satisfies these conditions is $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
$(3)\Rightarrow(1)$: If $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$, then $G[U]$ contains exactly one non-tree component $K_{3}$. Hence, $|\mathcal{N}_{1}|=1$.
This completes the proof of the claim. ∎
Based on the preceding analysis, we can now establish the following theorem, which characterizes the structure of the extremal graph $G$ that satisfies the property $P_{t,r}$ under the condition $43t_{0}\leq g_{0}<8t_{0}-$ 4.
Theorem 4.1.
Let $G$ be the spectral extremal graph that satisfies the property $P_{t,r}$. If $43t_{0}\leq g_{0}<8t_{0}-$ 4, then $G$ is isomorphic to the join of a clique $K_{\lfloor t\rfloor}$ and a graph $H$, where $H\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$. In other words,
$G\cong K_{\lfloor t\rfloor}\nabla\left(K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}\right).$
Now, we consider the case where $g_{0}$ does not satisfy the inequality $43t_{0}\leq g_{0}<8t_{0}-$ 4. By a similar argument as in Claim 4.3, if $\mathcal{N}_{1}$ contains a non-tree component $H$ that is not a $K_{3}$, then we can transform $H\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ into a star $K_{1,e(H)+a\lfloor\frac{1}{t_{0}}\rfloor-a}$, where $a=e(H)-|V(H)|+1$.
Through straightforward computation, we readily observe that $\eta(K_{1,e(H)+a\lfloor\frac{1}{t_{0}}\rfloor-a})\leq\eta(H)\leq g_{0}$. Given that there are enough $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ available, this transformation results in a new graph that satisfies the property $P_{t,r}$. By Lemma 4.3, the star $K_{1,e(H)+a\lfloor\frac{1}{t_{0}}\rfloor-a}$ is walk-preferable to $H\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$. Then, by Lemma 4.1, the transformed graph has a greater spectral radius, leading to a contradiction. Therefore, any non-tree component in $\mathcal{N}_{1}$ must be a $K_{3}$.
By Claim 4.5, $\mathcal{N}_{1}$ cannot contain any non-tree components at all. Consequently, this implies that $\mathcal{N}_{1}$ must be empty. Thus, $G[U]$ is a forest. By Claim 4.2, $G[U]$ consists solely of stars.
Following a similar discussion as in Claim 4.4, by using edge-shifting operations, we can make a star become as large as possible under the constraint of the property $P_{t,r}$. We then classify these stars based on their sizes.
Claim 4.6.
The stars in $G[U]$ can be categorized as follows:
  • Maximal star: $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$, where $a\geq 0$.
  • Big star: $K_{1,\lfloor\frac{1}{t_{0}}\rfloor}$.
  • Small star: $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$.
  • Minimal star: $K_{1,\gamma-1}$, where $1\leq\gamma\leq\lfloor\frac{1}{t_{0}}\rfloor$.
Furthermore, there is at most one maximal (or minimal) star.
These cases are determined by $t_{0}$, $g_{0}$, ensuring $G$ maximizes the spectral radius while satisfying $P_{t,r}$. Notice that when $a=0$, a maximal star reduces to a large star, and when $\gamma=\lfloor\frac{1}{t_{0}}\rfloor-1$, a minimal star transforms into a small star. In particular, when $\gamma=1$, it becomes an isolated vertex.
To maintain consistency with the previous analysis and introduce new notations, we define the following: Let $b$ denote the number of big stars in $G[U]$. Note that when $a=0$, the number of big stars is $b+1$. And let $\phi_{0}=\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor})=(1+\lfloor\frac{1}{t_{0}}\rfloor)t_{0}-1$. Clearly, $0<\phi_{0}\leq t_{0}$. We express $g_{0}$ in the form:
$g_{0}=\alpha_{0}\phi_{0}+\beta_{0},$
where $\alpha_{0}$ is a nonnegative integer, and $0\leq\beta_{0}<\phi_{0}$. Let $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor\frac{1}{t_{0}}\rfloor+\gamma_{0}$, where $1\leq\gamma_{0}\leq\lfloor\frac{1}{t_{0}}\rfloor$. For convenience, we denote $\eta^{+}(G[U])=\sum_{\eta(H)>0}\eta(H)$, where $H$ is a component of $G[U]$. Using these definitions, we now proceed to determine the number of edges in $G[U]$.
Claim 4.7.
$e(G[U])=n-\lfloor t\rfloor-(n_{0}+1)$.
Proof.
Since $G$ satisfies the property $P_{t,r}$, it holds that
$g_{0}\geq\eta^{+}(G[U])=at_{0}+(b+1)\phi_{0}=(a+b+1)\phi_{0}+a(1-t_{0}\lfloor\frac{1}{t_{0}}\rfloor)\geq(a+b+1)\phi_{0}.$
If $a+b+1\geq\alpha_{0}+1$, then we have
$g_{0}\geq\eta^{+}(G[U])\geq(a+b+1)\phi_{0}\geq\alpha_{0}\phi_{0}+\phi_{0}>\alpha_{0}\phi_{0}+\beta_{0}=g_{0},$
which is a contradiction. Therefore, $a+b+1\leq\alpha_{0}$.
Since $G[U]$ is a forest, the number of components is $\big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}$.
Thus, the number of edges in $G[U]$ is
$e(G[U])=n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}.$ (12)
Combining this with $a+b+1\leq\alpha_{0}$, we have
$e(G[U])\leq n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}=n-\lfloor t\rfloor-(n_{0}+1).$
Next, we show that
$e(G[U])\geq n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}.$
Suppose, for contradiction, that
$e(G[U])\leq n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}-1.$
Let $H$ be the union of the maximal star, minimal star, all big stars, and some small stars. By edge-shifting operations, we transform $H$ into a forest $H^{\prime}$, which is the union of $\alpha_{0}$ big stars and a new minimal star. Then, we construct a new graph $G^{\prime}$ with
$\eta^{+}(G^{\prime}[U])=\alpha_{0}\phi_{0}\leq g_{0}.$
This implies that $G^{\prime}$ satisfies the property $P_{t,r}$. Moreover,
$e(G^{\prime}[U])=n-\lfloor t\rfloor-\big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor\frac{1}{t_{0}}\rfloor}\big{\rceil}\geq e(G[U])+1,$
which implies $\rho(G^{\prime})>\rho(G)$, a contradiction. This completes the proof. ∎
Next, we determine the values of $a$ and $b$. To achieve this, we introduce new parameters. Let $a+b+c+1=\alpha_{0}$, where $c\geq 0$. Since $n-\lfloor t\rfloor=\alpha_{0}+n_{0}\lfloor\frac{1}{t_{0}}\rfloor+\gamma_{0}$, combining with equality (12), we have
$\gamma_{0}+\alpha_{0}-(a+b+1)\leq\lfloor\frac{1}{t_{0}}\rfloor.$
It follows that
$c\leq\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}.$
Let $G=G_{(a,b,c)}$, and denote $\eta^{+}(G_{(a,b,c)}[U])=\eta^{+}_{(a,b,c)}$. Clearly,
$\eta^{+}_{(0,\alpha_{0}-1,0)}=\alpha_{0}\phi_{0}=g_{0}-\beta_{0}.$
This implies that $G_{(0,\alpha_{0}-1,0)}$ satisfies the property $P_{t,r}$. Since $G_{(a,b,c)}$ is the spectral extremal graph, it also satisfies the property $P_{t,r}$. Therefore,
$\eta^{+}_{(a,b,c)}=(a+b+1)\phi_{0}+a(t_{0}-\phi_{0})=\alpha_{0}\phi_{0}+at_{0}-(a+c)\phi_{0}\leq g_{0}=\alpha_{0}\phi_{0}+\beta_{0}.$
This yields the following inequalities:
$at_{0}-(a+c)\phi_{0}$$\leq$ $\beta_{0},$ (13)
$0\leq c$$\leq$ $\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0},$ (14)
$a+c$$\leq$ $\alpha_{0}-1.$ (15)
The inequalities (13), (14), and (15) provide constraints on the parameters $a$, $b$, and $c$, which determine the structure of $G_{(a,b,c)}$. These constraints ensure that $G_{(a,b,c)}$ maximizes the spectral radius while satisfying the property $P_{t,r}$.
Notice that $G_{(a,b,c)}$ can be obtained from $G_{(0,\alpha_{0}-1,0)}$ by the following two types of edge-shifting operations:
  • (I)
    $H_{1}:=(a+1)K_{1,\lfloor\frac{1}{t_{0}}\rfloor}\implies H^{\prime}_{1}:=K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$.
  • (II)
    $H_{2}:=cK_{1,\lfloor\frac{1}{t_{0}}\rfloor}\cup K_{1,\gamma_{0}-1}\implies H^{\prime}_{2}:=cK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}\cup K_{1,c+\gamma_{0}-1}$.
We define $f_{k}(a,c)=W_{k}(H_{1}\cup H_{2})-W_{k}(H^{\prime}_{1}\cup H^{\prime}_{2})$. It is easy to see that $f_{1}(a,c)=0$. For $f_{2}(a,c)$, we have
$f_{2}(a,c)$ $=(a+\lfloor\frac{1}{t_{0}}\rfloor)^{2}+(c+\gamma_{0}-1)^{2}+(a+c)(\lfloor\frac{1}{t_{0}}\rfloor-1)^{2}$ (16)
$\quad-(a+c+1)\lfloor\frac{1}{t_{0}}\rfloor^{2}-(\gamma_{0}-1)^{2}$
$=a^{2}+2a\lfloor\frac{1}{t_{0}}\rfloor+c^{2}+2c(\gamma_{0}-1)-2(a+c)\lfloor\frac{1}{t_{0}}\rfloor+(a+c)$
$=a^{2}+a+c^{2}-c-2c\big{(}\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}\big{)}.$
We know that the spectral extremal graph is obtained precisely when the function $f(a,c)$ reaches its maximum. Therefore, our goal is to determine the values of $a$ and $c$ that maximize $f(a,c)$.
Claim 4.8.
Let $G_{(a,b,c)}$ be the spectral extremal graph. Then:
$a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}\quad\text{and}\quad c=0.$
Proof.
First, observe that the function $f_{2}(a,c)$ is convex with respect to $a$ and $c$. The symmetry plane of $f(a,c)$ with respect to $c$ is given by:
$c=\big{\lfloor}\frac{1}{t_{0}}\big{\rfloor}-\gamma_{0}+\frac{1}{2}.$
For $0\leq\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}<\lfloor\frac{1}{t_{0}}\rfloor-\gamma_{0}+\frac{1}{2}$, the extremum of $f_{2}(a,c)$ must occur at $c=0$.
Next, according to inequalities (13) and (15), we derive the following bounds for $a$ when $c=0$:
$a\leq\frac{\beta_{0}}{t_{0}-\phi_{0}}\quad\text{and}\quad a\leq\alpha_{0}-1.$
Since $a$ is a nonnegative integer and the symmetry plane of $f(a,c)$ with respect to $a$ is $a=-\frac{1}{2}$, it follows that
$a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}.$
Note that when $t_{0}=\phi_{0}$, we have $a=\alpha_{0}-1$.
Let $a^{*},b^{*},c^{*}$ be the values of $a,b,c$ that maximize $f_{2}(a,c)$. Then, the optimal values are:
$a^{*}=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}},\quad b^{*}=\alpha_{0}-1-a^{*},\quad c^{*}=0.$
Consequently, the spectral extremal graph $G_{(a^{*},b^{*},c^{*})}$ is obtained. ∎
The spectral extremal graph $G_{(a^{*},b^{*},c^{*})}$ is fully characterized by the optimal values $a^{*}$, $b^{*}$, and $c^{*}$, which are derived from the constraints and the optimization of $f_{2}(a,c)$. This graph maximizes the spectral radius while satisfying the property $P_{t,r}$. By determining the precise structure of $G_{(a^{*},b^{*},c^{*})}$, we complete the structural analysis of the extremal graph $G$.
Recall that $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor\frac{1}{t_{0}}\rfloor+\gamma_{0}$, where $1\leq\gamma_{0}\leq\lfloor\frac{1}{t_{0}}\rfloor$. Based on the above arguments, when $g_{0}$ does not satisfies that $43t_{0}\leq g_{0}<8t_{0}-$ 4, we have the following theorem.
Theorem 4.2.
Let $G$ be the spectral extremal graph that satisfies the property $P_{t,r}$. If $g_{0}$ does not satisfies that $43t_{0}\leq g_{0}<8t_{0}-$ 4, then $G$ is isomorphic to the join of a copy of $K_{\lfloor t\rfloor}$ and a graph $H$, where $H$ consists of stars as follows:
  • One maximal star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$, where $a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}$.
  • $b$ big stars $K_{1,\lfloor\frac{1}{t_{0}}\rfloor}$, where $b=\alpha_{0}-1-a$.
  • $\ell$ small stars $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$, where $\ell=n_{0}-b-1$.
  • One minimal star $K_{1,\gamma-1}$, where $\gamma=\gamma_{0}$.
Note that when $a=0$, it signifies that there is no maximal star in $G[U]$. Consequently, the number of big stars in $G[U]$ simplifies to $\alpha_{0}$. In the case $\gamma_{0}=\lfloor\frac{1}{t_{0}}\rfloor$, it indicates that there is no minimal star present in $G[U]$. The absence of a minimal star changes the composition of small stars. Specifically, the number of small stars becomes $n_{0}-b$.
In summary, by using Theorems 4.1 and 4.2, we have successfully completed the proof of Theorem 1.3. To conclude this section, we will present a corollary.
If $t$ is a positive integer, then $t_{0}=1$. Consequently, we have $\phi_{0}=1$, which implies $a=\alpha_{0}-1$ and $b=0$. Therefore, by applying Theorems 4.1 and 4.2, we obtain the following corollary:
Corollary 4.1.
If $t$ is a positive integer, then the spectral extremal graph $G$ is either $G\cong K_{\lfloor t\rfloor}\nabla\big{(}K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}\big{)}$ or $G\cong K_{\lfloor t\rfloor}\nabla\big{(}K_{1,\alpha_{0}}\cup(n-\lfloor t\rfloor-\alpha_{0}-1)K_{1}\big{)}$.
5 Concluding remarks
We may also consider a generalization of the Hereditarily Bounded Property, where the hereditary bound on the number of edges is not restricted to linear constraints. In his thesis, Guiduli posed the following problem:
Problem 5.1.
If for a graph $G$, $|E(H)|\leq c\cdot|V(H)|^{2}$ holds for all $H\subseteq G$, does it follow that $\rho(G)\leq 2c\cdot|V(G)|$? What can be said if the exponent of $2$ is replaced by some other constant less than $2$?
This issue has also been highlighted by Liu and Ning in their survey of unsolved problems in spectral graph theory [16]. It is worth noting that $\rho(G)\leq\sqrt{2c}n$ serves as a trivial upper bound, and that $2c$ would be best possible, as seen by Wilf’s theorem. If this question were true, then the spectral Erdős-Stone-Simonovits theorem [17] would follow from the Erdős-Stone-Simonovits theorem [9, 10] and Wilf’s theorem would follow from Turán’s theorem [19].