### v3

# Rui Lia,  Anyao Wangb,   Mingqing Zhaic

Abstract
We resolve a problem posed by Guiduli (1996) on the spectral radius of graphs satisfying the Hereditarily Bounded Property $P_{t,r}$, which requires that every subgraph $H$ with $|V(H)|\geq t$ satisfies $|E(H)|\leq t|V(H)|+r$. For an $n$-vertex graph $G$ satisfying $P_{t,r}$, where $t>0$ and $r\geq-\binom{\lfloor t+1\rfloor}{2}$, we prove that the spectral radius $\rho(G)$ is bounded above by $\rho(G)\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$, where $s=\binom{\lfloor t\rfloor+1}{2}+r$, thus affirmatively answering Guiduli’s conjecture.
Furthermore, we present a complete characterization of the extremal graphs that achieve this bound. These graphs are constructed as the join graph $K_{\lfloor t\rfloor}\nabla F$, where $F$ is either $K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$ or a forest consisting solely of star structures. The specific structure of such forests is meticulously characterized. Central to our analysis is the introduction of a novel potential function $\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|$, which quantifies the structural ”positivity” of subgraphs. By combining edge-shifting operations with spectral radius maximization principles, we establish sharp bounds on $\eta^{+}(G)$, the cumulative positivity of $G$. Our results contribute to the understanding of spectral extremal problems under edge-density constraints and provide a framework for analyzing similar hereditary properties.
Keywords: Hereditarily Bounded Property; spectral radius; extremal graph; edge-shifting
AMS   Classification: 05C50; 05C35
1 Introduction
All graphs considered in this context are simple, undirected and free from isolated vertices. We use $|G|$ and $e(G)$ to denote the numbers of vertices and edges in a graph $G$, respectively. For a subset $S$ of $V(G)$, let $G[S]$ be the subgraph induced by $S$. Let $N_{U}(v)$ denote the set of neighbors of vertex $v$ within the subset $U$ of vertices, formally defined as $N_{U}(v)=\{u\mid u\in U\text{ and }uv\in E(G)\}$. In particular, when $U=V(G)$, we simply write $N(v)$ for $N_{V(G)}(v)$. Denoted by $d_{U}(v)$ the cardinality of the set $N_{U}(v)$. Moreover, let $e(U)$ denote the size of $G[U]$ and $e(U,W)$ be the number of edges connecting $U$ and $W$ in $G$. Let $kG$ be the disjoint union of $k$ copies of $G$, and $G_{1}\nabla G_{2}$ be the graph obtained by joining each vertex of $G_{1}$ with each vertex of $G_{2}$.
Let $A(G)$ denote the adjacency matrix of a graph $G$, and let $\rho(G)$ represent the spectral radius of $A(G)$. A graph $G$ is said to have the Hereditarily Bounded Property $P_{t,r}$ if for each subgraph $H\subseteq G$ with $|V(H)|\geq t$, the size of $H$ satisfies $|E(H)|\leq t\cdot|V(H)|+r$, where $t>0$. In fact, many classes of graphs possess the Hereditarily Bounded Property, including planar graphs, which have been extensively studied in the context of spectral extremal problems (see [1, 3, 6, 7, 8, 15, 21]). Research into the spectral radius of graphs with similar hereditary properties has also flourished. For instance, constraints on matching numbers have been explored by Feng et al. [11], while constraints on cycle lengths were investigated by Gao et al. [12]. These studies contribute to a broader understanding of how the structural properties of graphs influence their spectral characteristics.
This concept was first introduced by Guiduli in his Ph.D. thesis [13], wherein he proved the following theorem.
Theorem 1.1.
Let $t\in\mathbf{N}$ and $r\geq-\binom{t+1}{2}$. If $G$ is a graph on $n$ vertices with the property $P_{t,r}$, then $\rho(G)\leq(t-1)/2+\sqrt{t(t+1)+2r}+\sqrt{tn}$.
For ease of notation, Guiduli defined $s=\binom{t+1}{2}+r$ in his thesis. When $t$ and $r$ are not restricted to integers, this result implies that $\rho(G)\leq c(s,t)+\sqrt{\lceil t\rceil n}$, where $c(s,t)$ represents a function of $s$ and $t$. Guiduli also posed the following problem.
Problem 1.1.
Let $t,r\in\mathbf{R}$, $t>0$ and $r\geq-\binom{\lfloor t+1\rfloor}{2}$. Is it true that $G$ has the property $P_{t,r}$, then $\rho(G)\leq c(s,t)+\sqrt{tn}$?
Furthermore, Guiduli conjectured a potentially stronger statement:
$\rho(G)\leq c(s,t)+\sqrt{\lfloor t\rfloor n}.$
In this paper, we affirmatively resolve the conjecture posed by Guiduli. Specifically, we establish the following theorem:
Theorem 1.2.
Let $t,r\in\mathbf{R}$, $t>0$ and $r\geq-\binom{\lfloor t+1\rfloor}{2}$. If $G$ is an $n$-vertex graph with the property $P_{t,r}$ , then $\rho(G)\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$.
Finally, we consider the structure of $G$ when it attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. The following theorem provides a precise characterization of such graphs.
Theorem 1.3.
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. Then $G$ is the join of two graphs: one is $K_{\lfloor t\rfloor}$, another is either $K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$ or a forest consisting of stars.
In Section 4, we provide a detailed characterization of the forest structure in Theorem 1.3, classifying the stars into maximal, big, small, and minimal stars based on their sizes. Therefore, all the extremal graphs are completely characterized.
2 Stability theorem
Let $\mathcal{G}$ denote the family of $n$-vertex graphs that satisfy property $P_{t,r}$. For any $G\in\mathcal{G}$, we assume for simplicity that the graph $G$ considered in this paper is connected. If $G$ were not connected, we could identify a component $C$ within $G$ such that the induced subgraph $G[C]$ also satisfies property $P_{t,r}$ and has the same spectral radius as $G$, i.e., $\rho(G[C])=\rho(G)$. Therefore, it suffices to consider only the connected component $G[C]$ for our analysis. Furthermore, we may assume that $n$ is sufficiently large. If $n$ were a finite constant, we could set $c(s,t)=n$. In this case, the spectral radius would be bounded by $\rho(G)\leq\rho(K_{n})=n-1\leq c(s,t)+\sqrt{\lfloor t\rfloor n}$. Hence, Theorem 1.2 would hold trivially.
Thus, for the remainder of our discussion, we focus on connected graphs $G\in\mathcal{G}$ with a sufficiently large order $n$. Given that $G\in\mathcal{G}$, it implies that $t\geq 1$. To see why, suppose for contradiction that $10<t<$ 1. For sufficiently large $n$, $e(G)\leq tn+r<n-1$. However, since $G$ is connected, $e(G)\geq n-1$. This leads to a contradiction, hence $t\geq 1$.
By the well-known Perron-Frobenius theorem, there exists a positive unit eigenvector $X$ corresponding to the spectral radius $\rho(G)$. Let $X=(x_{1},x_{2},\ldots,x_{n})^{T}$ and let $u^{*}\in V(G)$ such that $x_{u^{*}}=\max_{u\in V(G)}x_{u}.$ For simplicity, we will write $\rho=\rho(G)$.
For convenience, we define $s=\big{\lceil}\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\big{\rceil}+1$ in this paper. This definition of $s$ differs from that given in Guiduli’s thesis but similarly depends on $t$ and $r$. It is clear that $s>t+1$ as $r\geq-\binom{\lfloor t\rfloor+1}{2}$. To prove our main result, we introduce a stability theorem that plays a crucial role in our analysis.
Theorem 2.1.
Let $G$ be a graph of sufficiently large order $n$. Let $X$ be a non-negative eigenvector corresponding to $\rho(G)$, and $x_{u^{*}}=\max_{u\in V(G)}x_{u}$. If $G$ has the property $P_{t,r}$ and $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, then there exists a vertex subset $V$ of $G$ with $|V|=\lfloor t\rfloor$ such that $x_{v}\geq(1-\frac{1}{50s^{2}})x_{u^{*}}$ and $d_{G}(v)\geq(1-\frac{1}{25s^{2}})n$ for every $v\in V$.
Next, we will prove Theorem 2.1. We’ll define the set of vertices with large entries in $X$ relative to $x_{u^{*}}$, provide an upper bound for its cardinality, and then determine its exact value. This method was first introduced by Cioabă et al. to address a conjecture by Nikiforov [18], which comprises two parts: a spectral version of Erdős-Sós conjecture [4] and the spectral even cycle problem [5]. It was further developed by Byrne et al. [2].
Let $V_{1}=\{v\in V(G):\ x_{v}\geq(5s)^{-4}x_{u^{*}}\}$, $V_{2}=\{v\in V(G):\ x_{v}\geq(5s)^{-5}x_{u^{*}}\}$ and $V=\{v\in V(G):\ x_{v}\geq(5s)^{-1}x_{u^{*}}\}$. Obviously, $V\subseteq V_{1}\subseteq V_{2}$.
Claim 2.1.
$|V_{1}|\leq(5s)^{-6}n$ and $|V_{2}|\leq(5s)^{-5}n$.
Proof.
Let $v\in V_{1}$. We start with the inequality
$\rho x_{v}\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}(5s)^{-4}x_{u^{*}}.$
Moreover, given that $n$ is sufficiently large, it follows that $\rho x_{v}\geq 2s(5s)^{6}x_{u^{*}}.$ Additionally, we have $\rho x_{v}=\sum_{u\in N(v)}x_{u}\leq|N(v)|x_{u^{*}},$ where $N(v)$ denotes the set of neighbors of $v$ in $G$. Therefore,
$\sum_{v\in V_{1}}\rho x_{v}\leq\sum_{v\in V_{1}}|N(v)|x_{u^{*}}\leq\sum_{v\in V(G)}|N(v)|x_{u^{*}}=2e(G)x_{u^{*}}.$
Since $G\in\mathcal{G}$ and $s>t+1$, we know that $e(G)\leq tn+r\leq sn$. Combining this with the previous inequalities, we obtain
$2s(5s)^{6}|V_{1}|\leq\sum_{v\in V_{1}}\rho x_{v}\leq 2e(G)\leq 2sn.$
Consequently, $|V_{1}|\leq(5s)^{-6}n$, as desired. Similarly, $|V_{2}|\leq(5s)^{-5}n.$
For any vertex $u\in V(G)$, let $N_{2}(u)$ denote the set of all vertices that are at a distance of two from $u$ in the graph $G$. We can categorize the vertices around $u$ into several sets based on their distances from $u$ and their memberships in $V_{1}$ and $V_{2}$. Specifically, with respect to $V_{1}$, we denote
  • $A_{1}=V_{1}\cap N(u)$, $B_{1}=N(u)\setminus A_{1}$, $A_{2}=V_{1}\cap N_{2}(u)$, $B_{2}=N_{2}(u)\setminus A_{2}$.
Similarly, concerning the set $V_{2}$, we have
  • $C_{1}=V_{2}\cap N(u)$, $D_{1}=N(u)\setminus C_{1}$, $C_{2}=V_{2}\cap N_{2}(u)$, $D_{2}=N_{2}(u)\setminus C_{2}$.
Claim 2.2.
For every $u\in V(G)$, we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}\leq|N(u)|x_{u}+\Big{(}\frac{2sn}{(5s)^{6}}+\frac{2sn}{(5s)^{4}}\Big{)}x_{u^{*}}+\sum_{v\in B_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{w}.$ (1)
Proof.
Given that $\rho(G)\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$ and an arbitrary vertex $u\in V(G)$, we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}\leq\rho^{2}x_{u}=\sum_{v\in N(u)}\rho x_{v}=|N(u)|x_{u}+\sum_{v\in N(u),w\in N(v)\setminus\{u\}}x_{w}.$ (2)
Note that $N(u)=A_{1}\cup B_{1}$. Then $N(v)\setminus\{u\}=N_{A_{1}\cup A_{2}}(v)\cup N_{B_{1}\cup B_{2}}(v)$ for each $v\in N(u)$. By the definition of $V_{1}$, it follows that $x_{w}<(5s)^{-4}x_{u^{*}}$ for every $w\in B_{1}\cup B_{2}$. Therefore, we can establish an upper bound for the vertices in $A_{1}$ as follows:
$\sum_{v\in A_{1},w\in N(v)\setminus\{u\}}x_{w}$ $=\sum_{v\in A_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{u^{*}}+\sum_{v\in A_{1},w\in N_{B_{1}\cup B_{2}}(v)}(5s)^{-4}x_{u^{*}}$
$\leq\big{(}2e(A_{1})+e(A_{1},A_{2})\big{)}x_{u^{*}}+e(A_{1},B_{1}\cup B_{2})(5s)^{-4}x_{u^{*}}.$ (3)
Similarly, for vertices in $B_{1}$, we obtain
$\sum_{v\in B_{1},w\in N(v)\setminus\{u\}}x_{w}\leq\sum_{v\in B_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{w}+\big{(}2e(B_{1})+e(B_{1},B_{2})\big{)}(5s)^{-4}x_{u^{*}}.$ (4)
Since $G$ satisfies the property $P_{t,r}$ and $s>t+1$, we have that $2e(A_{1})+e(A_{1},A_{2})\leq 2s|V_{1}|$ and $e(A_{1},B_{1}\cup B_{2})+2e(B_{1})+e(B_{1},B_{2})\leq 2sn.$ Additionally, by Claim 2.1, we know that $|V_{1}|\leq(5s)^{-6}n$. Combining these three inequalities with our previous results (2-4), we can derive the desired inequality (1) directly. ∎
Using an argument analogous to that in Claim 2.2, with respect to $V_{2}$, we can find that for every $u\in V(G)$,
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}\leq|N(u)|x_{u}+\frac{4sn}{(5s)^{5}}x_{u^{*}}+\sum_{v\in D_{1},w\in N_{C_{1}\cup C_{2}}(v)}x_{w}.$ (5)
Claim 2.3.
$|V_{1}|<(5s)^{6}$.
Proof.
We begin by proving that $|N(u)|\geq(5s)^{-5}n$ for each $u\in V_{1}$. Suppose, for the sake of contradiction, that there exists a vertex $u_{0}\in V_{1}$ with $|N(u_{0})|<(5s)^{-5}n$.
By inequality (5), we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u_{0}}\leq|N(u_{0})|x_{u_{0}}+\frac{4sn}{(5s)^{5}}x_{u^{*}}+\sum_{v\in D_{1},w\in N_{C_{1}\cup C_{2}}(v)}x_{w}.$ (6)
Recall that $D_{1}\subseteq N(u_{0})$ and $C_{1}\cup C_{2}\subseteq V_{2}$. By the property $P_{t,r}$, we know that
$e(D_{1},C_{1}\cup C_{2})\leq e(N(u_{0})\cup V_{2})<s(|N(u_{0})|+|V_{2}|).$
Given our assumption, $|N(u_{0})|<(5s)^{-5}n$, and by Claim 2.1, $|V_{2}|\leq(5s)^{-5}n$. Hence,
$|N(u_{0})|x_{u_{0}}+\sum_{v\in D_{1},w\in N_{C_{1}\cup C_{2}}(v)}x_{w}\leq\Big{(}|N(u_{0})|+e(D_{1},C_{1}\cup C_{2})\Big{)}x_{u^{*}}<\frac{(2s+1)n}{(5s)^{5}}x_{u^{*}}.$
Combining this with inequality (6), we get that $\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u_{0}}<\frac{6s+1}{(5s)^{5}}nx_{u^{*}}.$ On the other hand, since $\lfloor t\rfloor\geq 1$ and $x_{u_{0}}\geq(5s)^{-4}x_{u^{*}}$ (because $u_{0}\in V_{1}$), we have
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u_{0}}\geq\frac{7}{10}nx_{u_{0}}>\frac{7s}{(5s)^{5}}nx_{u^{*}},$
leading to a contradiction. Therefore, $|N(u)|\geq(5s)^{-5}n$ for each vertex $u\in V_{1}$.
Summing this inequality over all vertices $u\in V_{1}$, we obtain
$|V_{1}|(5s)^{-5}n\leq\sum_{u\in V_{1}}|N(u)|\leq\sum_{u\in V(G)}|N(u)|=2e(G)<2sn.$
Thus, we conclude that $|V_{1}|<(5s)^{6}$. ∎
Claim 2.4.
For each $u\in V_{1}$, we have $|N(u)|\geq\big{(}\frac{x_{u}}{x_{u^{*}}}-\frac{1}{(5s)^{3}}\big{)}n$.
Proof.
Consider $S$ as a subset of vertices from $B_{1}$, with each vertex in $S$ having at least $\lfloor t\rfloor$ neighbors within $A_{1}\cup A_{2}$. Let $p=|A_{1}|+|A_{2}|$. We aim to prove that the size of $S$ is bounded by $|S|\leq s\cdot\ell$, where $\ell=\binom{p}{\lfloor t\rfloor}$.
If $p\leq\lfloor t\rfloor-1$, then it’s impossible for any vertex in $B_{1}$ to have $\lfloor t\rfloor$ neighbors in $A_{1}\cup A_{2}$, which implies that $S$ must be empty, and the statement holds trivially. Therefore, we focus on the case where $p\geq\lfloor t\rfloor$.
Suppose, for contradiction, that $|S|>s\cdot\ell$. Each vertex in $S$ connects to at least $\lfloor t\rfloor$ vertices in $A_{1}\cup A_{2}$. Considering all possible subsets of $A_{1}\cup A_{2}$ of size $\lfloor t\rfloor$, there are $\ell=\binom{p}{\lfloor t\rfloor}$ such subsets. Since each vertex in $S$ has at least $\lfloor t\rfloor$ neighbors, each of these vertices can potentially connect to any of the $\ell$ subsets.
Given that $|S|>s\cdot\ell$, this would mean that on average, each $\lfloor t\rfloor$-vertex subset of $A_{1}\cup A_{2}$ must connect to more than $s$ vertices in $S$. By the pigeonhole principle, there must exist at least one specific $\lfloor t\rfloor$-vertex subset of $A_{1}\cup A_{2}$ that connects to more than $s$ vertices in $S$, implying that this subset has over $s$ common neighbors in $S$. Note that $u\notin A_{1}\cup A_{2}$ and $S\subseteq B_{1}\subseteq N(u)$, then $G$ contains a bipartite subgraph isomorphic to $K_{s,\lfloor t\rfloor+1}$. Since $G$ satisfies the property $P_{t,r}$, we have the inequality
$e(K_{s,\lfloor t\rfloor+1})=s(\lfloor t\rfloor+1)\leq t(s+\lfloor t\rfloor+1)+r.$
Solving for $s$ in the inequality above yields that $s\leq\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t},$ which contradicts the definition of $s$. Therefore, it must be that $|S|\leq s\cdot\ell$.
Given that $p\leq|V_{1}|<(5s)^{6}$, it follows that $\ell$ and $|S|$ are constants. For sufficiently large $n$, we can establish the following bounds.
First, consider the number of edges between $S$ and $A_{1}\cup A_{2}$. We have
$e(S,A_{1}\cup A_{2})\leq|S|(|A_{1}|+|A_{2}|)\leq|S||V_{1}|\leq(5s)^{-6}n.$
On the other hand, by the definition of $S$, any vertex in $B_{1}\setminus S$ has fewer than $\lfloor t\rfloor$ neighbors in $A_{1}\cup A_{2}$. Therefore,
$e(B_{1}\setminus S,A_{1}\cup A_{2})\leq|B_{1}\setminus S|(\lfloor t\rfloor-1)\leq|N(u)|(\lfloor t\rfloor-1).$
Combining these two inequalities, we obtain that
$e(B_{1},A_{1}\cup A_{2})\leq|N(u)|(\lfloor t\rfloor-1)+(5s)^{-6}n.$
It follows that
$\sum_{v\in B_{1},w\in N_{A_{1}\cup A_{2}}(v)}x_{w}\leq e\big{(}B_{1},A_{1}\cup A_{2}\big{)}x_{u^{*}}\leq\Big{(}|N(u)|(\lfloor t\rfloor-1)+(5s)^{-6}n\Big{)}x_{u^{*}}.$ (7)
Note that $x_{u}\leq x_{u^{*}}$. Substituting inequality (7) into (1) gives that
$\lfloor t\rfloor(n-\lfloor t\rfloor)x_{u}$ $\leq\Big{(}|N(u)|+\frac{(2s+1)n}{(5s)^{6}}+\frac{2sn}{(5s)^{4}}+|N(u)|(\lfloor t\rfloor-1)\Big{)}x_{u^{*}}$
$\leq\lfloor t\rfloor\Big{(}|N(u)|+\frac{3sn}{(5s)^{4}}\Big{)}x_{u^{*}}.$
Rearranging terms yields that $|N(u)|\geq(n-\lfloor t\rfloor)\frac{x_{u}}{x_{u^{*}}}-\frac{3sn}{(5s)^{4}}.$ For sufficiently large $n$, since $\lfloor t\rfloor\frac{x_{u}}{x_{u^{*}}}\leq\lfloor t\rfloor<\frac{2sn}{(5s)^{4}}$, we conclude that $|N(u)|\geq(\frac{x_{u}}{x_{u^{*}}}-\frac{1}{(5s)^{3}})n,$ as desired. ∎
In the following, we give the proof of Theorem 2.1. Recall that $V=\{v\in V(G):\ x_{v}\geq(5s)^{-1}x_{u^{*}}\}$. By the definition of $V$, it is clear that $u^{*}\in V\subseteq V_{1}$. We will show that $V$ is the desired set in Theorem 2.1.
Proof of Theorem 2.1: We start by proving lower bounds on $x_{u}$ and $|N(u)|$ for each vertex $u\in V$. Suppose, for the sake of contradiction, that there exists $u_{0}\in V$ such that $x_{u_{0}}<(1-\frac{1}{50s^{2}})x_{u^{*}}.$
Since $u_{0}\in V$, we have $x_{u_{0}}\geq\frac{x_{u^{*}}}{5s}.$ By Claim 2.4, we know that
$|N(u^{*})|\geq\Big{(}1-\frac{1}{(5s)^{3}}\Big{)}n\quad\text{and}\quad|N(u_{0})|\geq\Big{(}\frac{1}{5s}-\frac{1}{(5s)^{3}}\Big{)}n.$
Define the following four sets: $A_{1}^{*}=V_{1}\cap N(u^{*})$, $B_{1}^{*}=N(u^{*})\setminus A_{1}^{*}$, $A_{2}^{*}=V_{1}\cap N_{2}(u^{*})$ and $B_{2}^{*}=N_{2}(u^{*})\setminus A_{2}^{*}$. From Claim 2.3, we have $|V_{1}|<(5s)^{6}<\frac{n}{(5s)^{3}}$ for sufficiently large $n$. Thus,
$|B_{1}^{*}|\geq|N(u^{*})|-|V_{1}|\geq\Big{(}1-\frac{2}{(5s)^{3}}\Big{)}n.$
Therefore,
$\big{|}B_{1}^{*}\cap N(u_{0})\big{|}\geq\big{|}B_{1}^{*}\big{|}+\big{|}N(u_{0})\big{|}-n\geq\Big{(}\frac{1}{5s}-\frac{3}{(5s)^{3}}\Big{)}n\geq\frac{22n}{125s}.$ (8)
Given (8), $u_{0}$ must have neighbors in $B_{1}^{*}$. Since $B_{1}^{*}\subseteq N(u^{*})$, $u_{0}$ is within distance two from $u^{*}$, implying $u_{0}\in N(u^{*})\cup N_{2}(u^{*})$. Recalling that $u_{0}\in V\subseteq V_{1}$, we conclude $u_{0}\in A_{1}^{*}\cup A_{2}^{*}$. Using inequality (1) with $u=u^{*}$, we get
$\lfloor t\rfloor(n\!-\!\lfloor t\rfloor)x_{u^{*}}$ $\!\leq\!\Big{(}|N(u^{*})|\!+\!\frac{2sn}{(5s)^{6}}\!+\!\frac{2sn}{(5s)^{4}}\!+\!e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\setminus\{u_{0}\}\big{)}\Big{)}x_{u^{*}}\!+\!e(B_{1}^{*},\{u_{0}\})x_{u_{0}}$
$\!\leq\!\Big{(}|N(u^{*})|\!+\!\frac{2sn}{(5s)^{6}}\!+\!\frac{2sn}{(5s)^{4}}\!+\!e\big{(}B_{1}^{*},A_{1}^{*}\cup A_{2}^{*}\big{)}\Big{)}x_{u^{*}}\!+\!e(B_{1}^{*},\{u_{0}\})(x_{u_{0}}\!-\!x_{u^{*}}),$
where $x_{u_{0}}-x_{u^{*}}<-\frac{x_{u^{*}}}{2(5s)^{2}}$ by our assumption.
From (7), we know that $e(B_{1}^{*},A_{1}^{*}\cup A_{2}^{*})\leq(\lfloor t\rfloor-1)|N(u^{*})|+\frac{n}{(5s)^{6}}.$ Furthermore, for large enough $n$, it easy to see $\lfloor t\rfloor^{2}\leq\frac{n}{(5s)^{6}}.$ This leads to
$\lfloor t\rfloor n$ $\leq\lfloor t\rfloor|N(u^{*})|+\frac{2sn}{(5s)^{4}}+\frac{(2s+2)n}{(5s)^{6}}-\frac{e(B_{1}^{*},\{u_{0}\})}{2(5s)^{2}}$
$<\lfloor t\rfloor n+\frac{2.2sn}{(5s)^{4}}-\frac{e(B_{1}^{*},\{u_{0}\})}{2(5s)^{2}},$
which implies that $e(B_{1}^{*},\{u_{0}\})<\frac{22n}{125s},$ contradicting (8). Therefore, $x_{u}\geq(1-\frac{1}{50s^{2}})x_{u^{*}}$ for each $u\in V$. By Claim 2.4, this also means
$|N(u)|\geq\Big{(}1-\frac{1}{2(5s)^{2}}-\frac{1}{(5s)^{3}}\Big{)}n\geq\big{(}1-\frac{1}{(5s)^{2}}\big{)}n.$
To prove that $|V|=\lfloor t\rfloor$, we begin by assuming that $|V|\geq\lfloor t\rfloor+1$. Based on the previous inequality, every vertex $u\in V$ has at most $\frac{n}{(5s)^{2}}$ non-neighbors. Hence, any $\lfloor t\rfloor+1$ vertices in $V$ have at least $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}$ common neighbors, where $n-(\lfloor t\rfloor+1)\frac{n}{(5s)^{2}}\geq s$ as $s>t+1$. Then $G$ contains a subgraph isomorphic to $K_{\lfloor t\rfloor+1,s}$. By a similar argument to that in the proof of Claim 2.4, there is a contradiction since $K_{\lfloor t\rfloor+1,s}$ does not satisfy the property $P_{t,r}$. Therefore, $|V|\leq\lfloor t\rfloor.$
Next, suppose for contradiction that $|V|\leq\lfloor t\rfloor-1.$ Since $u^{*}\in V\setminus(A_{1}^{*}\cup A_{2}^{*})$, we have $|V\cap(A_{1}^{*}\cup A_{2}^{*})|\leq\lfloor t\rfloor-2.$ This implies that
$e\big{(}B_{1}^{*},V\cap(A_{1}^{*}\cup A_{2}^{*})\big{)}\leq(\lfloor t\rfloor-2)n.$
On the other hand, considering the edges between $B_{1}^{*}$ and $(A_{1}^{*}\cup A_{2}^{*})\setminus V$, we have
$e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\setminus V\big{)}\leq e(G)\leq tn+r<sn.$
By the definition of $V$, we know that $x_{w}<\frac{x_{u^{*}}}{5s}$ for each $w\in(A_{1}^{*}\cup A_{2}^{*})\setminus V$. Now, setting $u=u^{*}$ in inequality (1), we obtain
$\lfloor t\rfloor(n\!-\!\lfloor t\rfloor)x_{u^{*}}$ $\leq\Big{(}|N(u^{*})|\!+\!\frac{2.5sn}{(5s)^{4}}\!+\!e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\cap V\big{)}\Big{)}x_{u^{*}}\!+\!e\big{(}B_{1}^{*},(A_{1}^{*}\cup A_{2}^{*})\setminus V\big{)}\frac{x_{u^{*}}}{5s}$
$\leq\Big{(}n\!+\!\frac{n}{5}\!+\!(\lfloor t\rfloor\!-\!2)n\!+\!\frac{n}{5}\Big{)}x_{u^{*}}$
$=\big{(}\lfloor t\rfloor\!-\!\frac{3}{5}\big{)}nx_{u^{*}}.$
Dividing both sides by $x_{u^{*}}$ and rearranging terms, we find that
$\lfloor t\rfloor(n-\lfloor t\rfloor)\leq\big{(}\lfloor t\rfloor-\frac{3}{5}\big{)}n.$
Simplifying further, we have $\lfloor t\rfloor^{2}\geq\frac{3}{5}n$, a contradiction. Therefore, $|V|=\lfloor t\rfloor$.
Combining both parts, we establish the desired results: $x_{u}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ and $|N(u)|\geq\big{(}1-\frac{1}{25s^{2}}\big{)}n$ for each vertex $u\in V$. Moreover, $|V|=\lfloor t\rfloor$. Thus, we complete the proof of Theorem 2.1. ∎
3 Proof of Theorem 1.2
Let $G$ be a graph that achieves the maximal spectral radius in $\mathcal{G}$. It is not hard to verify that $K_{\lfloor t\rfloor,n-\lfloor t\rfloor}\in\mathcal{G}$. Therefore, $\rho(G)\geq\rho(K_{\lfloor t\rfloor,n-\lfloor t\rfloor})=\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$.
Let $V\subseteq V(G)$ be a set defined in accordance with Theorem 2.1. We define $U$ as the set of vertices that are complete to $V$, meaning every vertex in $U$ is adjacent to every vertex in $V$. It is clear that $U$ and $V$ are disjoint sets, i.e., $U\cap V=\emptyset$. We then define $W=V(G)\setminus(U\cup V)$.
Claim 3.1.
$|U|\geq\frac{n}{2}$.
Proof.
Set $V=\{v_{1},v_{2},\dots,v_{\lfloor t\rfloor}\}$. By the Inclusion-Exclusion Principle, we have
$|U|=\big{|}\cap_{i=1}^{\lfloor t\rfloor}N(v_{i})\big{|}\geq\sum\limits_{i=1}^{\lfloor t\rfloor}\big{|}N(v_{i})\big{|}-(\lfloor t\rfloor-1)\big{|}\cup_{i=1}^{\lfloor t\rfloor}N(v_{i})\big{|}.$
It follows that
$|U|\geq\Big{(}1-\frac{1}{25s^{2}}\Big{)}n\cdot\lfloor t\rfloor-(\lfloor t\rfloor-1)n=\Big{(}1-\frac{\lfloor t\rfloor}{25s^{2}}\Big{)}n\geq\frac{n}{2},$
as claimed. ∎
Claim 3.2.
$W=\emptyset$.
Proof.
Choose a vertex $u\in U\cup W$. By Claim 3.1, let $N_{U}(u)=\{u_{1},u_{2},\dots,u_{d}\}$ and denote by $H$ the subgraph in $G$ induced by the set $V\cup N_{U}(u)\cup\{u\}$. Then $e(H)\geq(\lfloor t\rfloor+1)d$. Since $H$ has the property $P_{t,r}$, we have $e(H)\leq t|H|+r=t(\lfloor t\rfloor+1+d)+r$. Combining the two inequalities, we have $(\lfloor t\rfloor+1-t)d\leq t(\lfloor t\rfloor+1)+r$, and so
$d_{U}(u)=d\leq\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\leq s-1.$ (9)
for each vertex $u\in U\cup W.$
Note that $W$ is not complete to $V$, meaning each vertex in $W$ is not adjacent to some vertex in $V$. By Theorem 2.1, every vertex in $V$ is not adjacent to at most $\frac{n}{25s^{2}}$ vertices of $G$. Recall that $s>t+1$. Therefore,
$|W|\leq\frac{\lfloor t\rfloor n}{25s^{2}}\leq\frac{n}{25s}.$ (10)
Note that $d(u)\leq d_{U}(u)+|V|+d_{W}(u)\leq s-1+\lfloor t\rfloor+d_{W}(u)$. Moreover, since $G$ has the property $P_{t,r}$, either $e(W)\leq t|W|+r$ or $|W|<t$. Consequently, we get that
$\sum\limits_{w\in W}\rho x_{w}\leq\Big{(}\big{(}s-1+\lfloor t\rfloor\big{)}|W|+2e(W)\Big{)}x_{u^{*}}\leq\frac{(s+3t)\lfloor t\rfloor n}{25s^{2}}x_{u^{*}}.$
It follows that
$\sum\limits_{w\in N_{W}(v)}x_{w}\leq\sum\limits_{w\in W}x_{w}<\frac{(s+3t)\lfloor t\rfloor n}{25\rho s^{2}}x_{u^{*}}.$ (11)
Now for $u\in U\cup W$, we have $\rho x_{u}\leq(s-1+\lfloor t\rfloor)x_{u^{*}}+\sum_{w\in N_{W}(u)}x_{w}$. Combining with the inequality (11) and $\rho\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$, we deduce that
$x_{u}<\Big{(}\frac{s-1+\lfloor t\rfloor}{\rho}+\frac{(s+3t)\lfloor t\rfloor n}{25\rho^{2}s^{2}}\Big{)}x_{u^{*}}\leq\frac{s+4t}{25s^{2}}x_{u^{*}}<\frac{x_{u^{*}}}{5s}.$ (12)
Suppose that $W\neq\emptyset$. Let $G^{*}$ denote the graph obtained from $G$ by removing all edges within $W$ and from $W$ to $U$, then adding all possible edges from $W$ to $V$. Additionally, let $w_{1}$ be a vertex of minimal degree in $G[W]$, and let $E_{1}$ be the set of edges incident to $w_{1}$ in $G[W]$. Following the same approach, let $w_{2}$ be a vertex of minimal degree in $G[W\setminus\{w_{1}\}]$, and let $E_{2}$ be the set of edges incident to $w_{2}$ in $G[W\setminus\{w_{1}\}]$. Repeating the process, we define $W_{1}=W$ and $W_{i}=W\setminus\{w_{1},\dots,w_{i-1}\}$ for $i\geq 2$, where $w_{i}$ is a vertex of minimum degree in $G[W_{i}]$ and $E_{i}$ is the set of edges that are incident to $w_{i}$ in $G[W_{i}]$. By continuing this process iteratively, we have
$E(G[W])=\cup_{i=1}^{|W|}E_{i}=\cup_{i=1}^{|W|}\{w_{i}w:~{}w\in N_{W_{i}}(w)\}.$
Recall that $G\in\mathcal{G}$ and $s=\big{\lceil}\frac{t(\lfloor t\rfloor+1)+r}{\lfloor t\rfloor+1-t}\big{\rceil}+1\geq t^{2}+r+1$. When $|W_{i}|\geq t$, we have
$d_{W_{i}}(w_{i})\leq\frac{2e(W_{i})}{|W_{i}|}\leq\frac{2t|W_{i}|+2r}{|W_{i}|}\leq 2t+\frac{2r}{t}<2s.$
When $|W_{i}|<t$, we also get $d_{W_{i}}(w_{i})\leq\frac{2\binom{|W_{i}|}{2}}{|W_{i}|}<t<2s.$ Now, we can drive that
$\rho(G^{*})-\rho(G)$ $\geq X^{T}\big{(}A(G^{*})-A(G)\big{)}X$
$=\sum\limits_{i=1}^{|W|}2x_{w_{i}}\Big{(}\sum\limits_{v\in V\setminus N_{V}(w_{i})}x_{v}-\sum\limits_{u\in N_{U}(w_{i})}x_{u}-\sum\limits_{w\in N_{W_{i}}(w_{i})}x_{w}\Big{)}.$ (13)
In view of the definition of $W$, we have $|V\setminus N_{V}(w_{i})|\geq 1$ for each $w_{i}\in W$. By Theorem 2.1, $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for each $v\in V$. Moreover, by inequality (9), we know that $d_{U}(u)\leq s-1$ for each $u\in U$, and based on the previous discussion, we have $d_{W_{i}}(w_{i})<2s$ for each $w_{i}\in W_{i}$. Combining (12) and (3), we can see that
$\rho(G^{*})-\rho(G)\geq\sum\limits_{w_{i}\in W}2x_{w_{i}}\Big{(}\big{(}1-\frac{1}{50s^{2}}\big{)}-\big{(}3s-1\big{)}\frac{1}{5s}\Big{)}x_{u^{*}}>0.$
Based on the maximality of $\rho(G)$, the new graph $G^{*}$ cannot satisfy the property $P_{t,r}$. Specifically, there exists an induced subgraph $H$ of $G^{*}$ such that $|H|\geq t$ and $e(H)>t|H|+r$. Since $G^{*}[U\cup V]=G[U\cup V]$, the violation of the property $P_{t,r}$ must involve vertices in $W$, i.e., $V(H)\cap W\neq\emptyset$.
Let $H^{\prime}=G^{*}[V(H)\setminus W]$. According to the definition of $G^{*}$, we know that $N_{G^{*}}(w)=V$ for each $w\in W$. Hence, $e(H)=e(H^{\prime})+|V(H)\cap W|\cdot|V(H)\cap V|.$ Since $|V(H)\cap V|\leq\lfloor t\rfloor\leq t$ and $e(H)>t|H|+r$, it follows that
$e(H^{\prime})>t|V(H)|+r-t|V(H)\cap W|=t|V(H^{\prime})|+r.$
If $|V(H^{\prime})|\geq t$, the above inequality leads to a contradiction because $G^{*}[V(H^{\prime})]=G[V(H^{\prime})]$ and $G$ is assumed to satisfy the property $P_{t,r}$.
If $|V(H^{\prime})|<t$, then by Claim 3.1 and the inequality in (10), we have $|U\setminus V(H^{\prime})|\geq\frac{n}{2}-t\geq|V(H)\cap W|$. Since $U\setminus V(H)=U\setminus V(H^{\prime})$, we can choose a subset $U^{\prime}$ of $U\setminus V(H)$ with $|U^{\prime}|=|V(H)\cap W|$. Denote $H^{\prime\prime}=G^{*}[U^{\prime}\cup(V(H)\setminus W)]$. Then $|V(H^{\prime\prime})|=|V(H)|$, and $N_{G^{*}}(w)=V\subseteq N_{G^{*}}(u)$ for every $w\in W$ and every $u\in U^{\prime}$. Thus, $e(H^{\prime\prime})\geq e(H)>t|V(H)|+r=t|V(H^{\prime\prime})|+r$. This also leads to a contradiction.
In both cases, we arrive at a contradiction, which implies that our initial assumption $W\neq\emptyset$ must be false. Therefore, $W=\emptyset$. ∎
By Claim 3.2, we have $V(G)=V\cup U$ and $e(V,U)=|V||U|$, where $|V|=\lfloor t\rfloor$. Based on (9), every $u\in U$ has at most $s-1$ neighbors in $U$. Let $u^{\star}\in U$ such that $x_{u^{\star}}=\max_{u\in U}x_{u}$. Then $\rho x_{u^{\star}}\leq(s-1)x_{u^{\star}}+\lfloor t\rfloor x_{u^{*}}.$ It follows that $x_{u^{\star}}\leq\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}.$ Also, since $u^{*}\in V$, we have $\rho x_{u^{*}}\leq(\lfloor t\rfloor-1)x_{u^{*}}+(n-\lfloor t\rfloor)x_{u^{\star}}.$ Combining these two inequalities, we get that
$(\rho-\lfloor t\rfloor+1)x_{u^{*}}\leq(n-\lfloor t\rfloor)\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}.$
Dividing both sides by $x_{u^{*}}$, we obtain $\rho-\lfloor t\rfloor+1\leq\frac{(n-\lfloor t\rfloor)\lfloor t\rfloor}{\rho-s+1}.$ Expanding and rearranging terms, we have
$\rho^{2}-(s+\lfloor t\rfloor-2)\rho-\lfloor t\rfloor n+\lfloor t\rfloor^{2}+(s-1)(\lfloor t\rfloor-1)\leq 0.$
Solving this quadratic inequality, we find that
$\rho\leq\frac{(s+\lfloor t\rfloor-2)}{2}+\sqrt{\frac{(s+\lfloor t\rfloor-2)^{2}}{4}+\lfloor t\rfloor n-\lfloor t\rfloor^{2}-(s-1)(\lfloor t\rfloor-1)}.$
Therefore, $\rho\leq c(s,t)+\sqrt{\lfloor t\rfloor n},$ where $c(s,t)$ is a constant depending on $s$ and $t$. This completes the proof of Theorem 1.2.
4 Characterization of extremal graphs
In this section, we will present the proof Theorem 1.3. Before delving into our proof, we introduce some concepts and several lemmas. An $\ell$-walk consists of $\ell+1$ vertices where consecutive vertices are adjacent. Let $w_{G}^{\ell}(u)$ be the number of $\ell$-walks starting at vertex $u$ in $G$. The total number of $\ell$-walks in $G$ is defined as $W_{\ell}(G)=\sum_{v\in V(G)}w_{G}^{\ell}(v)$. Obviously, $W_{1}(G)=2e(G)$ and $W_{2}(G)=\sum_{v\in V(G)}(d_{G}(v))^{2}.$
Let $G_{1},G_{2}$ be two graphs, potentially of different orders. We say that $G_{1}$ is walk-equivalent to $G_{2}$ and denote this as $G_{1}\equiv G_{2}$, if $W_{\ell}(G_{1})=W_{\ell}(G_{2})$ for all $\ell\geq 1$. If there exists an $\ell$ such that $W_{\ell}(G_{1})>W_{\ell}(G_{2})$ and $W_{i}(G_{1})=W_{i}(G_{2})$ for $i\leq\ell-1$, we say that $G_{1}$ is strictly walk-preferable to $G_{2}$ and denote this as $G_{1}\succ G_{2}$.
The following theorem, provided by Zhang, offers a framework for comparing the spectral radius of graphs based on their local walk structures. This seems particularly useful in various graph-theoretic analyses and applications.
Lemma 4.1.
([22]) Let $G_{0}$ be an $n$-vertex connected graph. Assume that $S$ is a non-empty proper subset of $V(G_{0})$ and $T$ is a set of isolated vertices in $G_{0}-S$ such that $e(S,T)=|S||T|$. For $i\in\{1,2\}$, let $H_{i}$ be a graph with vertex set $T$, and let $G_{i}$ be the graph obtained from $G_{0}$ by embedding the edges of $H_{i}$ into $T$. When $\rho(G_{0})$ is sufficiently large compared to $|T|$, we have the following conclusions:
(i) If $H_{1}\equiv H_{2}$, then $\rho(G_{1})=\rho(G_{2})$.
(ii) If $H_{1}\succ H_{2}$, then $\rho(G_{1})>\rho(G_{2})$.
Let $\mathbb{G}(n,m)$ be the family of graphs with $n$ vertices and $m$ edges, and let $F_{p}(H)$ be the sum of the $p$-th powers of vertex degrees of a graph $H$, formally defined as
$F_{p}(H)=\sum_{v\in V(H)}\big{(}d_{H}(v)\big{)}^{p},$
where $p\geq 1$ is an integer. Ismailescu and Stefanica showed the following result regarding the maximization of $F_{p}(H)$.
Lemma 4.2.
([14]) Let $H^{*}\in\mathbb{G}(n,m)$ such that $F_{p}(H^{*})=\max_{H\in\mathbb{G}(n,m)}F_{p}(H)$, where $p\geq 2$ is an integer. If $n\geq m+2$, then $H^{*}$ is the star $K_{1,m}$ plus $n-m-1$ isolated vertices. Moreover, it is unique except in the case when $p=2$ and $m=3$ (when both $K_{1,3}\cup(n-4)K_{1}$ and $K_{3}\cup(n-3)K_{1}$ are extremal graphs).
Given that $p=2$ and $n=m-1$, and following a similar approach to Lemma 4.2, the result below can be derived.
Lemma 4.3.
Let $H^{*}\in\mathbb{G}(n,m)$ such that $F_{2}(H^{*})=\max_{H\in\mathbb{G}(n,m)}F_{2}(H)$. If $m=n-1$, then $H^{*}\cong K_{1,m}$. Moreover, this configuration is unique unless $m=3$, in which case both $K_{1,3}$ and $K_{3}\cup K_{1}$ are extremal graphs.
Given a graph $F$, we define $\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|$, and let $\mathcal{C}(F)$ be the set of all components of $F$. It is easy to see that $\eta(F)=\sum_{F_{i}\in\mathcal{C}(F)}\eta(F_{i})$. Furthermore, a component $F_{i}$ of $F$ is said to be positive if $\eta(F_{i})>0$. Let $\mathcal{C}^{+}(F)$ be the set of all positive components of $F$. We denote
$\eta^{+}(F)=\sum_{F_{i}\in\mathcal{C}^{+}(F)}\eta(F_{i}),\ t_{0}=\lfloor t\rfloor+1-t\ \text{and}\ g_{0}=t\lfloor t\rfloor+r-\binom{\lfloor t\rfloor}{2}.$
Specially, if $\mathcal{C}^{+}(F)=\emptyset$, then $\eta^{+}(F)=0$. Recall that $r\geq-\binom{\lfloor t\rfloor+1}{2}$. Clearly, $t_{0}$ and $g_{0}$ are two constants with $0<t_{0}\leq 1$ and $g_{0}\geq 0$. Based on the definitions of $g_{0}$ and $\eta(F)$, we have the following two lemmas.
Lemma 4.4.
If $F^{\prime}$ is a subgraph of $F$, then $\eta(F^{\prime})\leq\eta(F)$.
Proof.
We first assume that $F$ is connected. If $V(F^{\prime})=V(F)$, then it is clear that $\eta(F^{\prime})\leq\eta(F)$. Next, suppose that $V(F^{\prime})\subsetneq V(F)$, and let $F_{1}^{\prime},\dots,F_{p}^{\prime}$ be the components obtained by removing $V(F^{\prime})$ from $F$. Since each $F_{i}^{\prime}$ is connected, $e(F_{i}^{\prime})\geq|V(F_{i}^{\prime})|-1$, and there is at least one edge connecting $F_{i}^{\prime}$ to $F^{\prime}$. Thus
$\eta(F)$ $=e(F)+(\lfloor t\rfloor-t)|V(F)|$
$\geq e(F^{\prime})+\sum_{i=1}^{p}e(F_{i}^{\prime})+\sum_{i=1}^{p}e(F_{i}^{\prime},F^{\prime})+(\lfloor t\rfloor-t)|V(F^{\prime})|+(\lfloor t\rfloor-t)\sum_{i=1}^{p}|V(F_{i}^{\prime})|.$
Combining $e(F_{i}^{\prime})\geq|V(F_{i}^{\prime})|-1$ and $e(F_{i}^{\prime},F^{\prime})\geq 1$, we have
$\eta(F)\geq\eta(F^{\prime})+\sum_{i=1}^{p}(|V(F_{i}^{\prime})|-1)+k+(\lfloor t\rfloor-t)\sum_{i=1}^{p}|V(F_{i}^{\prime})|=\eta(F^{\prime})+t_{0}\sum_{i=1}^{p}|V(F_{i}^{\prime})|,$
where $t_{0}=\lfloor t\rfloor+1-t>0$. It follows that $\eta(F^{\prime})\leq\eta(F)$.
Now we consider the case when $F$ is not connected. Let $\mathcal{C}(F)=\{F_{1},\dots,F_{q}\}$. Based on the above argument, we know that $\eta\big{(}F^{\prime}[V(F^{\prime})\cap V(F_{i})]\big{)}\leq\eta(F_{i})$ for $i\in\{1,\ldots,q\}$. It follows that $\eta(F^{\prime})\leq\sum_{i=1}^{q}\eta(F_{i})=\eta(F)$, as desired. ∎
Lemma 4.5.
$K_{\lfloor t\rfloor}\nabla F$ satisfies the property $P_{t,r}$ if and only if $\eta^{+}(F)\leq g_{0}$.
Proof.
First, assume that $K_{\lfloor t\rfloor}\nabla F$ satisfies the property $P_{t,r}$. Recall that $g_{0}=t\lfloor t\rfloor+r-\binom{\lfloor t\rfloor}{2}\geq 0$. If $\mathcal{C}^{+}(F)=\emptyset$, then $\eta^{+}(F)=0\leq g_{0}$, and the conclusion holds trivially. Otherwise, let $F^{+}$ be the union of all positive components of $F$. Since $K_{\lfloor t\rfloor}\nabla F^{+}$ is an induced subgraph of $K_{\lfloor t\rfloor}\nabla F$ with $|V(K_{\lfloor t\rfloor}\nabla F^{+})|\geq t$, it must satisfy $P_{t,r}$, and hence,
$e(K_{\lfloor t\rfloor}\nabla F^{+})=e(F^{+})+\lfloor t\rfloor|V(F^{+})|+\binom{\lfloor t\rfloor}{2}\leq t\big{(}\lfloor t\rfloor+|V(F^{+})|\big{)}+r.$
Rearranging this inequality, we obtain that $e(F^{+})+(\lfloor t\rfloor-t)|V(F^{+})|\leq g_{0},$ which implies that $\eta^{+}(F)=\eta(F^{+})\leq g_{0}$.
Now, assume that $\eta^{+}(F)\leq g_{0}$. We will prove that $K_{\lfloor t\rfloor}\nabla F$ satisfies the property $P_{t,r}$. Let $H$ be an arbitrary subgraph of $K_{\lfloor t\rfloor}\nabla F$ with $|V(H)|\geq t$. It suffices to show that $e(H)\leq t|V(H)|+r.$ Denote $F^{\prime}=F[V(H)\cap V(F)]$ and $t^{\prime}=|V(H)\cap V(K_{\lfloor t\rfloor})|$, where $0\leq t^{\prime}\leq\lfloor t\rfloor$. By Lemma 4.4, we have $\eta(F^{\prime})\leq\eta(F)\leq\eta^{+}(F)\leq g_{0},$ and hence
$e(F^{\prime})=\eta(F^{\prime})-(\lfloor t\rfloor-t)|V(F^{\prime})|\leq g_{0}-(\lfloor t\rfloor-t)(|V(H)|-t^{\prime}).$ (14)
Observe that $e(H)\leq e(F^{\prime})+\binom{t^{\prime}}{2}+t^{\prime}(|V(H)|-t^{\prime})$, and recall that $g_{0}=t\lfloor t\rfloor+r-\binom{\lfloor t\rfloor}{2}$. By combining (14), we obtain that
$e(H)\leq t|V(H)|+r-(\lfloor t\rfloor-t^{\prime})(|V(H)|-t^{\prime}-t)+\binom{t^{\prime}}{2}-\binom{\lfloor t\rfloor}{2}.$
Note that $\lfloor t\rfloor\geq t^{\prime}$ and $|V(H)|\geq t$. Define $f(t^{\prime})=(\lfloor t\rfloor-t^{\prime})t^{\prime}+\binom{t^{\prime}}{2}-\binom{\lfloor t\rfloor}{2}$. Then
$e(H)\leq t|V(H)|+r+f(t^{\prime}).$
If $t^{\prime}=\lfloor t\rfloor$, then $f(t^{\prime})=0$ and thus $e(H)\leq t|V(H)|+r$, as required. If $t^{\prime}\leq\lfloor t\rfloor-1$, it is easy to check that $f(\lfloor t\rfloor-1)=0$ and $f^{\prime}(t^{\prime})=\lfloor t\rfloor-t^{\prime}-\frac{1}{2}>0$. Thus $f(t^{\prime})\leq 0$ for $10\leq t^{\prime}\leq\lfloor t\rfloor-$ 1. Hence, $e(H)\leq t|V(H)|+r$. This completes the proof. ∎
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. By Claim 3.2, we know that $V(G)=V\cup U$, where $|V|=\lfloor t\rfloor$ and $e(V,U)=|V||U|$. Let $T$ be a subset of $U$ such that $e(T,U\setminus T)=0$ and $|T|$ is constant. Define two new graphs $H_{1}$ and $H_{2}$ with $V(H_{1})=V(H_{2})=T$ and $e(H_{1})=e(H_{2})$. Let $G_{i}$ be the graph obtained from $G$ by embedding the edges of $H_{i}$ into $T$ for $i\in\{1,2\}$. By Lemma 4.1, we obtain the following corollary.
Corollary 4.1.
If $H_{1}\succ H_{2}$ (specially, $F_{2}(H_{1})>F_{2}(H_{2})$), then $\rho(G_{1})>\rho(G_{2})$.
Let $g(V)=t\lfloor t\rfloor+r-e(V)$. It is obvious that $g_{0}\leq g(V)\leq t\lfloor t\rfloor+r$. Observe that $G[V]\nabla F$ is a spanning subgraph of $K_{\lfloor t\rfloor}\nabla F$. By a similar argument as in the first part of Lemma 4.5, we obtain the following statement.
Corollary 4.2.
If $G[V]\nabla F$ satisfies the property $P_{t,r}$, then $\eta^{+}(F)\leq g(V).$
We will now formally present the proof of Theorem 1.3, starting with the following key claim. The proof proceeds via a minimal counterexample.
Claim 4.1.
$V$ is a clique.
Proof.
Suppose to the contrary that $G[V]$ has a non-edge $v_{1}v_{2}$. Clearly, $\rho(G+v_{1}v_{2})>\rho(G)$. Then $G$ admits a subgraph $H$ of order at least $t$ such that $H+v_{1}v_{2}$ no longer satisfies the property $P_{t,r}$. Since $H$ satisfies the property $P_{t,r}$ but $H+v_{1}v_{2}$ no longer does, we conclude that $e(H)=\big{\lfloor}t|V(H)|+r\big{\rfloor}$. Now, $H+v_{1}v_{2}$ is called a counterexample of $G+v_{1}v_{2}$. We may assume that $H+v_{1}v_{2}$ is minimal with respect to the order.
Let $A=V\cap V(H)$, $B=U\cap V(H)$ and $C=V\setminus V(H)$. Define $L$ as the vertex set of all components in $G[U\setminus B]$ that have at least one edge connecting to some vertex in $B$. Furthermore, let $N_{0}$ denote the vertex set of all tree-components in $G[U\setminus(B\cup L)]$ and $N_{1}$ be the vertex set of all other components in $G[U\setminus(B\cup L)]$.
First, we will prove that the number of vertices in $L\cup N_{1}$ is finite. Consider the size of the induced subgraph $G[V\cup B\cup L\cup N_{1}]$:
$e(V\cup B\cup L\cup N_{1})$ $=e(V\cup B)+e(L\cup N_{1})+e(V\cup B,L\cup N_{1})$
$\geq e(H)+e(L)+e(N_{1})+e(B,L)+e(V,L\cup N_{1}).$
Recall that $e(H)=\big{\lfloor}t|V(H)|+r\big{\rfloor}$. Moreover, we see that $e(L)+e(B,L)\geq|L|$, $e(N_{1})\geq|N_{1}|$ and $e(V,L\cup N_{1})=\lfloor t\rfloor(|L|+|N_{1}|)$. Thus
$e(V\cup B\cup L\cup N_{1})\geq\big{\lfloor}t|V(H)|+r\big{\rfloor}+|L|+|N_{1}|+\big{\lfloor}t\big{\rfloor}\big{(}|L|+|N_{1}|\big{)}.$
On the other hand, by the property $P_{t,r}$, we have
$e(V\cup B\cup L\cup N_{1})\leq t(|V|+|B|+|L|+|N_{1}|)+r.$
Note that $|V|+|B|=|V(H)|+|C|$. Combining the above two inequalities, we obtain
$\big{(}\lfloor t\rfloor+1-t\big{)}\big{(}|L|+|N_{1}|\big{)}$ $\leq t\big{(}|V(H)|+|C|\big{)}+r-\big{\lfloor}t|V(H)|+r\big{\rfloor}$
$\leq t|C|+1\leq t\lfloor t\rfloor+1,$
which shows that $|L|+|N_{1}|$ is finite.
Recall that $t_{0}=\lfloor t\rfloor+1-t$. We now partition $N_{0}$ into two parts: Let $N_{0}^{1}$ denote the vertex set of all components in $G[N_{0}]$ with cardinality at least $\lfloor 1/{t_{0}}\rfloor+1$, and let $N_{0}^{2}$ denote the vertex set of all remaining components in $G[N_{0}]$.
Next, we will show that $|N_{0}^{1}|$ is also finite. Let $F$ be a component of $G[N_{0}^{1}]$. Note that $F$ is a tree with $|V(F)|\geq\lfloor 1/{t_{0}}\rfloor+1>1/{t_{0}}$, and recall that $\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|$. It follows that $\eta(F)=t_{0}|V(F)|-1>0$, and thus, $\eta^{+}(N_{0}^{1})=\eta(N_{0}^{1})=t_{0}|N_{0}^{1}|-c$, where $c$ is the number of the components in $G[N_{0}^{1}]$. Notice that $c(\lfloor 1/{t_{0}}\rfloor+1)\leq|N_{0}^{1}|$. Let $t^{\prime}_{0}=1/(\lfloor 1/{t_{0}}\rfloor+1)$. Then $t^{\prime}_{0}<t_{0}$, $c\leq t^{\prime}_{0}|N_{0}^{1}|$ and $\eta^{+}(N_{0}^{1})\geq(t_{0}-t_{0}^{\prime})|N_{0}^{1}|$. Since $G[V]\nabla G[N_{0}^{1}]$ satisfies the property $P_{t,r}$, by Corollary 4.2, we get that $\eta^{+}(N_{0}^{1})\leq g(V)\leq t\lfloor t\rfloor+r$. By combining these two inequalities, the assertion follows.
Thirdly, we shall demonstrate that $|B|$ is finite provided that $C\neq\emptyset$. Consider the subgraph $G[V(H)\cup C]$. It is easy to see that $e(V(H)\cup C)\geq e(H)+|C||B|$, where $e(H)=\big{\lfloor}t|V(H)|+r\big{\rfloor}\geq t|V(H)|+r-1$. However, using the property $P_{t,r}$, we have $e(V(H)\cup C)\leq t(|V(H)|+|C|)+r.$ Hence, $|C||B|\leq t|C|+1,$ and so $|B|\leq t+1$.
Observe that $e(B,N_{1}\cup N_{0}^{1}\cup N_{0}^{2})=0$. In view of (9), we have $e(B,L)\leq|L|(s-1)$. Building on the previous argument, we further see that both $e(B,L)$ and $e(L\cup N_{1}\cup N_{0}^{1})$ are constants. We proceed with the proof by considering two cases.
The first case is when $e(B)$ is finite. Now, $e(U\setminus N_{0}^{2})$ is also finite. Let $G^{\prime}$ be the graph obtained from $G$ by removing all edges within $U\setminus N_{0}^{2}$ and adding the edge $v_{1}v_{2}$. We assert that $G^{\prime}$ satisfies the property $P_{t,r}$. To verify this, observe that $E(G^{\prime}[U])=E(G^{\prime}[N_{0}^{2}])$. Let $F^{\prime}$ be a component of $G^{\prime}[U]$. Then $F^{\prime}$ is a tree with $|V(F^{\prime})|\leq\lfloor\frac{1}{t_{0}}\rfloor$. Thus
$\eta(F^{\prime})=e(F^{\prime})+(\lfloor t\rfloor-t)|V(F^{\prime})|=t_{0}|V(F^{\prime})|-1\leq 0,$ (15)
which implies that $G^{\prime}[U]$ contains no any positive components. Hence, $\eta^{+}(G^{\prime}[U])=0\leq g_{0}$. By Lemma 4.5, $K_{\lfloor t\rfloor}\nabla G^{\prime}[U]$ satisfies the property $P_{t,r}$. Since $G^{\prime}$ is a spanning subgraph of $K_{\lfloor t\rfloor}\nabla G^{\prime}[U]$, it follows that $G^{\prime}$ also satisfies the property $P_{t,r}$.
By Theorem 2.1, $x_{v}\geq\big{(}1-\frac{1}{50s^{2}}\big{)}x_{u^{*}}$ for $v\in V$. In view of (9), we have $x_{u}\leq\frac{\lfloor t\rfloor}{\rho-s+1}x_{u^{*}}$ for each $u\in U$. Now we see that
$\rho(G^{\prime})-\rho(G)\geq X^{T}\big{(}A(G^{\prime})-A(G)\big{)}X=2x_{v_{1}}x_{v_{2}}-\!\!\sum_{u_{i}u_{j}\in E(G[U\setminus N_{0}^{2}])}\!\!2x_{u_{i}}x_{u_{j}}.$
Recall that $\rho\geq\sqrt{\lfloor t\rfloor(n-\lfloor t\rfloor)}$ and $e(G[U\setminus N_{0}^{2}])$ is finite. Therefore, we can deduce that $\rho(G^{\prime})>\rho(G)$, which contradicts the maximality of $\rho(G)$.
It remains the case when $e(B)$ is sufficiently large. Now $B$ is an infinite set, and by the previous proof, we have $C=\emptyset$. Thus $V(H)=V\cup B$. We construct $G^{\prime\prime}$ as follows: (i) delete all edges in $G[L\cup N_{1}\cup N_{0}^{1}]$; (ii) delete all edges from $L$ to $B$; (iii) delete an edge in $G[B]$ and add the edge $v_{1}v_{2}$ to $G[V]$. Using a similar argument to the one for $G^{\prime}$, we have $\rho(G^{\prime\prime})>\rho(G)$, which implies that $G^{\prime\prime}$ does not satisfy the property $P_{t,r}$.
We conclude the proof by contradiction, demonstrating that $G^{\prime\prime}$ satisfies the property $P_{t,r}$. Observe that $e(G^{\prime\prime}[V\cup B])=e(G[V\cup B])$. Since $G$ satisfies the property $P_{t,r}$, we have $e(G^{\prime\prime}[V\cup B])=e(G[V\cup B])\leq t(|V|+|B|)+r$. Recall that $V(H)=V\cup B$ and $H+v_{1}v_{2}$ is a minimal counterexample. We see that $e(H^{\prime})\leq t|V(H^{\prime})|+r$ for any subgraph $H^{\prime}$ of $G^{\prime\prime}[V\cup B]$. Hence, $G^{\prime\prime}[V\cup B]$ satisfies the property $P_{t,r}$.
On the other hand, observe that $G^{\prime\prime}[U\setminus B]\cong G^{\prime}[U\setminus B]$, and every component $F^{\prime}$ of $G^{\prime\prime}[U\setminus B]$ is a tree with $|V(F^{\prime})|\leq\lfloor\frac{1}{t_{0}}\rfloor$. In view of (15), we know that $\eta(F^{\prime})\leq 0$ for each component $F^{\prime}$ of $G^{\prime\prime}[U\setminus B]$. It follows that $\eta(G^{\prime\prime}[U\setminus B])\leq 0$.
Let $H^{\prime\prime}$ be an arbitrary subgraph of $G^{\prime\prime}$. Moreover, let $H_{1}^{\prime\prime}=G^{\prime\prime}[V(H^{\prime\prime})\cap(V\cup B)]$ and $H_{2}^{\prime\prime}=G^{\prime\prime}[V(H^{\prime\prime})\cap(U\setminus B)]$. Since $G^{\prime\prime}[V\cup B]$ satisfies the property $P_{t,r}$, we have $e(H_{1}^{\prime\prime})\leq t|V(H_{1}^{\prime\prime})|+r$. By Lemma 4.4, $\eta(H_{2}^{\prime\prime})\leq\eta(G^{\prime\prime}[U\setminus B])\leq 0$, which yields that $e(H_{2}^{\prime\prime})+\lfloor t\rfloor|V(H_{2}^{\prime\prime})|\leq t|V(H_{2}^{\prime\prime})|.$ Note that there are no edges between $B$ and $U\setminus B$ in $G^{\prime\prime}$. Thus, $e(H^{\prime\prime})\leq e(H_{1}^{\prime\prime})+e(H_{2}^{\prime\prime})+\lfloor t\rfloor|V(H_{2}^{\prime\prime})|\leq t|V(H^{\prime\prime})|+r.$ Therefore, $G^{\prime\prime}$ satisfies the property $P_{t,r}$. This completes the proof. ∎
By Lemma 4.5 and Claim 4.1, we have $\eta^{+}(G[U])\leq g_{0}$. Next, we classify the components of $G[U]$ into three distinct categories $\mathcal{N}_{1},$ $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$, where
(i) $\mathcal{N}_{1}$ includes all components that are not trees;
(ii) $\mathcal{N}_{2}$ contains all tree components of order at least $\lfloor 1/t_{0}\rfloor+1$;
(iii) $\mathcal{N}_{3}$ consists of all tree components of order at most $\lfloor 1/t_{0}\rfloor$.
Claim 4.2.
Every component in $\mathcal{N}_{1}$ is positive. Moreover, $\sum_{F\in\mathcal{N}_{1}}|V(F)|\leq g_{0}/t_{0}.$
Proof.
For each $F\in\mathcal{N}_{1}$, we have $e(F)\geq|V(F)|$. Recall that $t_{0}=\lfloor t\rfloor+1-t>0$. Thus,
$\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|\geq t_{0}|V(F)|>0.$ (16)
Hence, $F$ is a positive component. Furthermore, summing over all components $F\in\mathcal{N}_{1}$, we obtain that $\sum_{F\in\mathcal{N}_{1}}\eta(F)\geq t_{0}\sum_{F\in\mathcal{N}_{1}}|V(F)|$. However, by Lemma 4.5, we know that $\sum_{F\in\mathcal{N}_{1}}\eta(F)\leq\eta^{+}(G[U])\leq g_{0}$ since $G\cong K_{\lfloor t\rfloor}\nabla G[U]$ and $G$ satisfies the property $P_{t,r}$. Therefore, we have $\sum_{F\in\mathcal{N}_{1}}|V(F)|\leq g_{0}/t_{0},$ as claimed. ∎
Claim 4.3.
Every component in $\mathcal{N}_{2}$ is positive and $\sum_{F\in\mathcal{N}_{2}}|V(F)|$ is constant.
Proof.
For each $F\in\mathcal{N}_{2}$, we have $e(F)=|V(F)|-1$ and $|V(F)|\geq\lfloor 1/t_{0}\rfloor+1$. Thus,
$\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|=t_{0}|V(F)|-1>0.$ (17)
Hence, all components in $\mathcal{N}_{2}$ are positive and $\sum_{F\in\mathcal{N}_{2}}\eta(F)=t_{0}\sum_{F\in\mathcal{N}_{2}}|V(F)|-|\mathcal{N}_{2}|$, where $\sum_{F\in\mathcal{N}_{2}}|V(F)|\geq|\mathcal{N}_{2}|(\lfloor 1/t_{0}\rfloor+1)$. Thus,
$\sum_{F\in\mathcal{N}_{2}}\eta(F)\geq\Big{(}t_{0}-\frac{1}{\lfloor 1/t_{0}\rfloor+1}\Big{)}\sum_{F\in\mathcal{N}_{2}}|V(F)|.$
On the other hand, by Lemma 4.5, we have $\sum_{F\in\mathcal{N}_{2}}\eta(F)\leq\eta^{+}(G[U])\leq g_{0}$. Therefore, $\sum_{F\in\mathcal{N}_{2}}|V(F)|$ is upper bounded by $g_{0}/\big{(}t_{0}-\frac{1}{\lfloor 1/t_{0}\rfloor+1}\big{)}.$
Claim 4.4.
Every component in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ is a star.
Proof.
Suppose to the contrary that $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ contains some components that are not stars. Then, let $G^{\prime}$ be the graph obtained from $G$ by replacing each component in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ with a star of the same order. By Lemma 4.3, the star is strictly walk-preferable to other trees with the same order. By Claim 4.3 and the definition of $\mathcal{N}_{3}$, every component in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ has finite order. Hence, by Corollary 4.1, we have $\rho(G^{\prime})>\rho(G)$.
However, in view of (17), we know that $\eta(F)=t_{0}|V(F)|-1$ for each tree $F\in\mathcal{N}_{2}\cup\mathcal{N}_{3}$. This implies that $\eta^{+}(G^{\prime}[U])=\eta^{+}(G[U])\leq g_{0}$. By Lemma 4.5, $G^{\prime}$ satisfies the property $P_{t,r}$, which leads to a contradiction. ∎
From (15), it follows that $\eta(F)\leq 0$ for every tree $F$ with $|V(F)|\leq\lfloor 1/t_{0}\rfloor$. Hence, no component in $\mathcal{N}_{3}$ is positive. Furthermore, the following statement holds for $\mathcal{N}_{3}$.
Claim 4.5.
There is at most one star $F$ in $\mathcal{N}_{3}$ with $|V(F)|\leq\lfloor 1/t_{0}\rfloor-1$.
Proof.
Suppose, for contradiction, that there are two stars $K_{1,a}$ and $K_{1,b}$ in $\mathcal{N}_{3}$ such that $20\leq b\leq a\leq\lfloor 1/t_{0}\rfloor-$ 2. If $b\geq 1$, we construct a graph $G^{\prime}$ from $G$ by performing an edge-shifting operation, i.e., replacing $K_{1,a}\cup K_{1,b}$ with $K_{1,a+1}\cup K_{1,b-1}$. In light of (15), neither $K_{1,a+1}$ nor $K_{1,b-1}$ are positive. Thus, $\eta^{+}(G^{\prime}[U])=\eta^{+}(G[U])\leq g_{0}$. By Lemma 4.5, $G^{\prime}$ satisfies the property $P_{t,r}$. Notice that $F_{2}(K_{1,a+1}\cup K_{1,b-1})>F_{2}(K_{1,a}\cup K_{1,b})$. By Corollary 4.1, we have $\rho(G^{\prime})>\rho(G)$, which contradicts the maximality of $\rho(G)$.
If $b=0$, i.e., $K_{1,b}$ is an isolated vertex, we construct a new graph $G^{\prime\prime}$ from $G$ by adding an edge between the center of $K_{1,a}$ and the isolated vertex. It is clear that $\rho(G^{\prime\prime})>\rho(G)$. Moreover, $\eta(K_{1,a+1})\leq 0$, and hence $\eta^{+}(G^{\prime\prime}[U])=\eta^{+}(G[U])\leq g_{0}$. Therefore, $G^{\prime\prime}$ satisfies the property $P_{t,r}$, which also leads to a contradiction. ∎
If $\mathcal{N}_{2}$ contains two stars of order at least $\lfloor 1/t_{0}\rfloor+2$, we can also perform an edge-shifting operation, i.e., replacing $K_{1,\lfloor 1/t_{0}\rfloor+a}\cup K_{1,\lfloor 1/t_{0}\rfloor+b}$ with $K_{1,\lfloor 1/t_{0}\rfloor+a+1}\cup K_{1,\lfloor 1/t_{0}\rfloor+b-1}$. Using a similar argument as in Claim 4.5, we obtain the following statement.
Claim 4.6.
There is at most one star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$ in $\mathcal{N}_{2}$ with $a\geq 1$.
By Claims 4.2 and 4.3, $\sum_{F\in\mathcal{N}_{3}}|V(F)|=\Theta(n)$. Furthermore, by Claim 4.5, $\mathcal{N}_{3}$ contains $\Theta(n)$ copies of $K_{1,\lfloor 1/{t_{0}}\rfloor-1}$. Based on this, we have the following assertion.
Claim 4.7.
There is at most one component $F$ in $\mathcal{N}_{1}$. If $|\mathcal{N}_{1}|=1$, then $F$ is a triangle.
Proof.
Suppose to the contrary that there are two components $F_{1},F_{2}\in\mathcal{N}_{1}$. Then $e(F_{i})\geq|V(F_{i})|\geq 3$ for $i\in\{1,2\}$. Let $\sum_{i=1}^{2}e(F_{i})=\sum_{i=1}^{2}|V(F_{i})|-1+a$. By Claim 4.2, we obtain $\sum_{i=1}^{2}|V(F_{i})|\leq g_{0}/t_{0}$, and thus $a$ is also bounded by a constant. Recall that $\mathcal{N}_{3}$ contains $\Theta(n)$ stars of order $\lfloor 1/t_{0}\rfloor$. Now, consider the subgraph $F$ formed by the union of $F_{1}$, $F_{2}$, and $a$ stars $K_{1,\lfloor 1/t_{0}\rfloor-1}$ from $\mathcal{N}_{3}$. It is clear that $e(F)=|V(F)|-1$, and $|V(F)|$ is finite. In view of (15), we have $\eta(K_{1,\lfloor 1/t_{0}\rfloor-1})\leq 0$. Recall that $t_{0}=1+\lfloor t\rfloor-t$. Then
$\eta^{+}(F)=\sum_{i=1}^{2}\eta(F_{i})=\sum_{i=1}^{2}e(F_{i})+(\lfloor t\rfloor-t)\sum_{i=1}^{2}|V(F_{i})|=t_{0}\sum_{i=1}^{2}|V(F_{i})|-1+a.$ (18)
Now, we transform $F$ into a star $F^{\prime}=K_{1,|V(F)|-1}$ and construct a graph $G^{\prime}$ from $G$ by replacing the edges of $F$ with the edges of $F^{\prime}$. From (17), we observe that $\eta^{+}(F^{\prime})=\eta(F^{\prime})=t_{0}|V(F^{\prime})|-1$, where $|V(F^{\prime})|=\sum_{i=1}^{2}|V(F_{i})|+a\lfloor{1}/{t_{0}}\rfloor$. By comparing with (18), it is straightforward to verify that $\eta^{+}(F^{\prime})\leq\eta^{+}(F)$. Thus, $\eta^{+}(G^{\prime}[U])\leq\eta^{+}(G[U])\leq g_{0}$, and hence $G^{\prime}$ satisfies the property $P_{t,r}$. Notice that $e(F^{\prime})=e(F)>3$. By Lemma 4.3, we see that $F_{2}(F^{\prime})>F_{2}(F)$. Since $|V(F)|$ is finite, by Corollary 4.1, we obtain $\rho(G^{\prime})>\rho(G)$, which leads to a contradiction. Therefore, $|\mathcal{N}_{1}|\leq 1$.
Next, assume that $|\mathcal{N}_{1}|=\{F_{1}\}$. We will demonstrate that $F_{1}$ is a triangle. The proof follows similarly to the previous one. If $e(F_{1})\geq 4$, we denote $e(F_{1})=|V(F_{1})|-1+a$ and $F=F_{1}\cup aK_{1,\lfloor 1/t_{0}\rfloor-1}$. Let $G^{\prime}$ be a graph obtained from $G$ by removing the edges of $F$ and adding the edges of $F^{\prime}$, where $F^{\prime}=K_{1,|V(F)|-1}$. We also have $\eta^{+}(F^{\prime})\leq\eta^{+}(F),$ and hence $\eta^{+}(G^{\prime}[U])\leq\eta^{+}(G[U])\leq g_{0}$, which implies that $G^{\prime}$ satisfies the property $P_{t,r}$. By a similar argument as above, we can conclude that $\rho(G^{\prime})>\rho(G)$, a contradiction. Therefore, $e(F_{1})=3$, and so $F_{1}$ is a triangle. This completes the proof. ∎
With the constraints on $\mathcal{N}_{1},\mathcal{N}_{2}$ and $\mathcal{N}_{3}$ established as outlined above, we can now relate these findings to the overall structure of $G[U]$. Specifically, we will investigate the implications of having a single component in $\mathcal{N}_{1}$.
Claim 4.8.
The following statements are equivalent:
(i) $|\mathcal{N}_{1}|=1$;
(ii) $43t_{0}\leq g_{0}<8t_{0}-$ 4;
(iii) $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
Proof.
We prove the equivalence of the three statements by showing the implications $(i)\Rightarrow(ii)$, $(ii)\Rightarrow(iii)$, and $(iii)\Rightarrow(i)$.
$(i)\Rightarrow(ii)$: Assume that $|\mathcal{N}_{1}|=1$. By Claim 4.7, the unique component in $\mathcal{N}_{1}$ is a triangle $K_{3}$. Recall that $t_{0}=\lfloor t\rfloor+1-t>0$. Then $\eta(K_{3})=e(K_{3})+(\lfloor t\rfloor-t)|V(K_{3})|=3t_{0}$. Since $\eta(K_{3})\leq\eta^{+}(G[U])\leq g_{0}$, we have $g_{0}\geq 3t_{0}$.
Next, we show that $g_{0}<8t_{0}-4$. Suppose, for contradiction, that $g_{0}\geq 8t_{0}-4$. Since $\mathcal{N}_{1}$ contains a triangle, we have $e(G[U])\geq 4$. Otherwise, $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ would consist only of isolated vertices. Recall that $\mathcal{N}_{3}$ contains $\Theta(n)$ components. We can construct a new graph $G^{\prime}$ from $G$ by replacing $K_{3}\cup 5K_{1}$ with $4K_{2}$. We observe that $\eta^{+}(G^{\prime}[U])\leq\eta(4K_{2})=8t_{0}-4\leq g_{0}$. Thus, by Lemma 4.5, $G^{\prime}$ satisfies the property $P_{t,r}$. However, since $4K_{2}$ is walk-preferable to $K_{3}\cup 5K_{1}$, Corollary 4.1 implies that $\rho(G^{\prime})>\rho(G)$, a contradiction. This contradiction establishes the existence of a star $K_{1,a}$ in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ with $a\geq 1$.
Now, we construct another graph $G^{\prime\prime}$ from $G$ by replacing $K_{3}\cup K_{1,a}$ with $K_{1,3+a}$. By Lemma 4.3 and Corollary 4.1, we have $\rho(G^{\prime\prime})>\rho(G)$. Since $K_{1,3+a}$ and $K_{3}\cup K_{1,a}$ have the same number of vertices and the same number of edges, we have $\eta(K_{1,3+a})=\eta(K_{3}\cup K_{1,a})$, and hence, $\eta^{+}(G^{\prime\prime}[U])=\eta^{+}(G[U])\leq g_{0}.$ Thus, $G^{\prime\prime}$ also satisfies the property $P_{t,r}$, contradicting the maximality of $\rho(G)$. Therefore, $g_{0}<8t_{0}-4$.
$(ii)\Rightarrow(iii)$: Assume that $43t_{0}\leq g_{0}<8t_{0}-$ 4. Then $t_{0}>\frac{4}{5}$. If $e(G[U])\leq 2$, we modify the graph $G$ by removing the edges within $U$ and then embedding a triangle in $U$. By a similar argument as above, we get a contradiction. Thus, $e(G[U])\geq 3$.
Let $F$ be the subgraph induced by all the edges in $G[U]$. Clearly, $|V(F)|\leq 2e(F)$, and thus, $\eta(F)=e(F)-(1-t_{0})|V(F)|\geq e(F)(2t_{0}-1).$ If $e(F)\geq 4$, then $\eta(F)\geq 8t_{0}-4$ and so $\eta^{+}(G[U])=\eta(F)>g_{0}$. This is a contradiction. Thus, we have $e(F)=3$. Furthermore, by Lemma 4.3, $F$ must be either $K_{1,3}$ or $K_{3}$. Since $W_{3}(K_{3})=24>W_{3}(K_{1,3})=18$, $K_{3}$ is walk-preferable to $K_{1,3}$. Therefore, $F\cong K_{3}$, and so $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
$(iii)\Rightarrow(i)$: If $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$, then $G[U]$ contains exactly one non-tree component $K_{3}$. Hence, $|\mathcal{N}_{1}|=1$.
This completes the proof. ∎
Building on the preceding analysis and Claim 4.8, we can now establish the following theorem, which uniquely determines the extremal graph $G$ when $43t_{0}\leq g_{0}<8t_{0}-$ 4.
Theorem 4.1.
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. If $43t_{0}\leq g_{0}<8t_{0}-$ 4, then we have
$G\cong K_{\lfloor t\rfloor}\nabla\left(K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}\right).$
In the following, we need only consider the case where $g_{0}$ does not satisfy the inequality $43t_{0}\leq g_{0}<8t_{0}-$ 4. By Claims 4.7 and 4.8, $\mathcal{N}_{1}$ must be empty. Thus, $G[U]$ is a forest. Furthermore, $G[U]$ consists solely of stars by Claim 4.4.
According to the definition of $\mathcal{N}_{2}$, the stars within $\mathcal{N}_{2}$ are of the form $K_{1,\lfloor 1/t_{0}\rfloor+a},$ where $a\geq 0$. By Claim 4.6, there are only two distinct types of stars in $\mathcal{N}_{2}$. We refer to the star as a big star when $a=0$, and as a maximal star for cases where $a\geq 1$. For $\mathcal{N}_{3}$, the stars are of the form $K_{1,\gamma-1}$, where $1\leq\gamma\leq\lfloor 1/t_{0}\rfloor$. These stars are referred to as small stars when $\gamma=\lfloor 1/t_{0}\rfloor$, and as minimal stars when $1\leq\gamma<\lfloor 1/t_{0}\rfloor$. By Claims 4.5 and 4.6, there exists at most one minimal (or maximal) star. In particular, when $\gamma=1$, a minimal star becomes an isolated vertex.
To determine the number of edges in $G[U]$, we introduce the following notations: Let $b$ denote the number of big stars (i.e., $K_{1,\lfloor 1/t_{0}\rfloor}$) in $G[U]$. For convenience, when considering the maximal star $K_{1,\lfloor 1/t_{0}\rfloor+a}$, allow $a=0$. If $a=0$, then the number of big stars is $b+1$; otherwise, there is exactly one maximal star of the form $K_{1,\lfloor 1/t_{0}\rfloor+a}$ with $a\geq 1$.
Define $\phi_{0}=\eta(K_{1,\lfloor 1/t_{0}\rfloor})=(1+\lfloor 1/t_{0}\rfloor)t_{0}-1$, noting that $0<\phi_{0}\leq t_{0}$. Express $g_{0}$ as
$g_{0}=\alpha_{0}\phi_{0}+\beta_{0},$ (19)
where $\alpha_{0}$ is a nonnegative integer and $0\leq\beta_{0}<\phi_{0}$.
Additionally, let $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor 1/t_{0}\rfloor+\gamma_{0}$, where $n_{0}$ is an integer sufficiently large, and $1\leq\gamma_{0}\leq\lfloor 1/t_{0}\rfloor$. Using these definitions, we now proceed to determine $e(G[U])$.
Claim 4.9.
$e(G[U])=n-\lfloor t\rfloor-(n_{0}+1)$.
Proof.
By Claim 4.3, the stars within $\mathcal{N}_{2}$ are positive. Moreover, recall that $\mathcal{N}_{3}$ contains no positive stars. Then, $\eta^{+}(G[U])=\eta(K_{1,\lfloor 1/t_{0}\rfloor+a})+b\eta(K_{1,\lfloor 1/t_{0}\rfloor}).$ It follows that
$\eta^{+}(G[U])=at_{0}+(b+1)\phi_{0}=(a+b+1)\phi_{0}+a(t_{0}-\phi_{0})\geq(a+b+1)\phi_{0}.$ (20)
Note that $\eta^{+}(G[U])\leq g_{0}$. If $a+b+1\geq\alpha_{0}+1$, then $g_{0}\geq(\alpha_{0}+1)\phi_{0}>\alpha_{0}\phi_{0}+\beta_{0}=g_{0}$, which is a contradiction. Therefore,
$a+b+1\leq\alpha_{0}.$ (21)
Since $G[U]$ is a forest, and based on the classification of $\mathcal{N}_{2}\cup\mathcal{N}_{3}$, the number of components in $G[U]$ is $\big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor 1/t_{0}\rfloor}\big{\rceil}$. Thus, the number of edges in $G[U]$ is
$e(G[U])=n-\lfloor t\rfloor-\Big{\lceil}\frac{n-\lfloor t\rfloor-(a+b+1)}{\lfloor 1/t_{0}\rfloor}\Big{\rceil}.$ (22)
Combining (21) and (22), we obtain that
$e(G[U])\leq n-\lfloor t\rfloor-\Big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor 1/t_{0}\rfloor}\Big{\rceil}=n-\lfloor t\rfloor-(n_{0}+1).$
Next, we show that $e(G[U])\geq n-\lfloor t\rfloor-(n_{0}+1)$. Suppose, to the contrary, that this is not the case. Let $F$ be the union of the maximal star, the minimal star, all big stars, and some (finite) small stars. By edge-shifting operations, we can transform $F$ into a forest $F^{\prime}$, which is the union of $\alpha_{0}$ big stars and a new minimal star. Then construct a graph $G^{\prime}$ from $G$ by replacing the edges of $F$ with the edges of $F^{\prime}$. It is clear that $\eta^{+}(G^{\prime}[U])=\alpha_{0}\phi_{0}\leq g_{0}$, which implies that $G^{\prime}$ satisfies the property $P_{t,r}$. Moreover,
$e(G^{\prime}[U])=n-\lfloor t\rfloor-\Big{\lceil}\frac{n-\lfloor t\rfloor-\alpha_{0}}{\lfloor 1/t_{0}\rfloor}\Big{\rceil}=n-\lfloor t\rfloor-(n_{0}+1)\geq e(G[U])+1.$
By Corollary 4.1, we have $\rho(G^{\prime})>\rho(G)$, a contradiction. This completes the proof. ∎
Next, we determine the values of $a$ and $b$. To achieve this, we shall introduce a new parameter. Denote $c=\alpha_{0}-(a+b+1)$. Based on (21), it follows that $c\geq 0$. Recall that $n-\lfloor t\rfloor=\alpha_{0}+\gamma_{0}+n_{0}\lfloor 1/t_{0}\rfloor$. Combining this with equality (22) and Claim 4.9, we have $\alpha_{0}+\gamma_{0}-(a+b+1)\leq\lfloor 1/t_{0}\rfloor$. It follows that $c\leq\lfloor 1/t_{0}\rfloor-\gamma_{0}$.
We can observe that $G$ is uniquely determined by $a$ and $b$. Therefore, we may denote $G=G_{(a,b,c)}$, and let $\eta^{+}(G_{(a,b,c)}[U])=\eta^{+}_{(a,b,c)}$. Recall that $\eta^{+}_{(a,b,c)}=\eta^{+}(G[U])\leq g_{0}$. Furthermore, based on (20), we obtain that
$\eta^{+}_{(a,b,c)}=at_{0}+(b+1)\phi_{0}=at_{0}+(\alpha_{0}-a-c)\phi_{0}\leq g_{0}=\alpha_{0}\phi_{0}+\beta_{0}.$ (23)
From the previous discuss, we get the following inequalities:
$at_{0}-(a+c)\phi_{0}$$\leq$ $\beta_{0},$ (24)
$0\leq c$$\leq$ $\lfloor 1/t_{0}\rfloor-\gamma_{0},$ (25)
$a+c$$\leq$ $\alpha_{0}-1.$ (26)
In view of (23), we have $\eta^{+}_{(0,\alpha_{0}-1,0)}=\alpha_{0}\phi_{0}=g_{0}-\beta_{0},$ and by Lemma 4.5, we see that $G_{(0,\alpha_{0}-1,0)}$ satisfies the property $P_{t,r}$. Observe that $G_{(0,\alpha_{0}-1,0)}$ contains $\alpha_{0}$ components of the form $K_{1,\lfloor 1/t_{0}\rfloor}$. Additionally, recall that $n-\lfloor t\rfloor-n_{0}\lfloor 1/t_{0}\rfloor=\alpha_{0}+\gamma_{0}$. This further implies that $G_{(0,\alpha_{0}-1,0)}$ contains a minimal component $K_{1,\gamma_{0}-1}$ and $n_{0}-\alpha_{0}$ components of the form $K_{1,\lfloor 1/t_{0}\rfloor-1}$. Recall that $\alpha_{0}=a+b+c+1$. It is easy to see that $G_{(a,b,c)}$ can be obtained from $G_{(0,\alpha_{0}-1,0)}$ by the following two types of edge-shifting operations:
(i) $F_{1}:=(a+1)K_{1,\lfloor 1/t_{0}\rfloor}\implies F^{\prime}_{1}:=K_{1,\lfloor 1/t_{0}\rfloor+a}\cup aK_{1,\lfloor 1/t_{0}\rfloor-1}$;
(ii) $F_{2}:=cK_{1,\lfloor 1/t_{0}\rfloor}\cup K_{1,\gamma_{0}-1}\implies F^{\prime}_{2}:=cK_{1,\lfloor 1/t_{0}\rfloor-1}\cup K_{1,c+\gamma_{0}-1}$.
We now define $f_{k}(a,c)=W_{k}(F^{\prime}_{1}\cup F^{\prime}_{2})-W_{k}(F_{1}\cup F_{2})$. Then, it is clear that $f_{1}(a,c)=0$ since $e(F^{\prime}_{1}\cup F^{\prime}_{2})=e(F_{1}\cup F_{2})$. Note that $W_{2}(F)=\sum_{v\in V(F)}\big{(}d_{F}(v)\big{)}^{2}$. We further have
$f_{2}(a,c)$ $=(a+\lfloor 1/t_{0}\rfloor)^{2}+(c+\gamma_{0}-1)^{2}+(a+c)(\lfloor 1/t_{0}\rfloor-1)^{2}$
$\quad-(a+c+1)\lfloor 1/t_{0}\rfloor^{2}-(\gamma_{0}-1)^{2}$
$=a^{2}+2a\lfloor 1/t_{0}\rfloor+c^{2}+2c(\gamma_{0}-1)-2(a+c)\lfloor 1/t_{0}\rfloor+(a+c)$
$=a^{2}+a+c^{2}-2c\big{(}\lfloor 1/t_{0}\rfloor-\gamma_{0}+1/2\big{)}.$
Subject to the property $P_{t,r}$, the extremal graph $G$ is obtained precisely when the function $f_{2}(a,c)$ reaches its maximum. Therefore, our goal is to determine the values of $a$ and $c$ that maximize $f_{2}(a,c)$, under the condition that $\eta^{+}_{(a,b,c)}\leq g_{0}$.
Claim 4.10.
$a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}$, $b=\alpha_{0}-a-1$ and $c=0$.
Proof.
First, observe that $f_{2}(a,c)$ is a quadratic function with respect to $a$ and $c$, and it represents an upward-opening paraboloid of revolution. The symmetry plane of $f(a,c)$ with respect to $c$ is $c=\big{\lfloor}1/t_{0}\big{\rfloor}-\gamma_{0}+\frac{1}{2}$. Form (25), we know that $0\leq c\leq\lfloor 1/t_{0}\rfloor-\gamma_{0}$. Then, the extremum of $f_{2}(a,c)$ must occur when $c=0$.
Next, from (24) and (26), we can derive that $a\leq\frac{\beta_{0}}{t_{0}-\phi_{0}}$ and $a\leq\alpha_{0}-1$ when $c=0$. Since $a$ is a nonnegative integer and the symmetry plane of $f_{2}(a,c)$ with respect to $a$ is $a=-\frac{1}{2}$, it follows that $a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}},$ where $a=\alpha_{0}-1$ when $t_{0}=\phi_{0}$. Recall that $c=\alpha_{0}-(a+b+1)$. We further have $b=\alpha_{0}-a-1$.
Now it suffices to verify that $\eta^{+}_{(a,b,c)}\leq g_{0}$ for the values of $a,b,c$ obtained above. In view of (23), we have
$\eta^{+}_{(a,b,c)}=at_{0}+(b+1)\phi_{0}=\alpha_{0}\phi_{0}+a(t_{0}-\phi_{0})\leq\alpha_{0}\phi_{0}+\beta_{0}=g_{0}.$
This completes the proof. ∎
Recall that $n-\lfloor t\rfloor-\alpha_{0}=n_{0}\lfloor 1/t_{0}\rfloor+\gamma_{0}$, where $1\leq\gamma_{0}\leq\lfloor 1/t_{0}\rfloor$. Based on the above arguments, the following theorem holds.
Theorem 4.2.
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. If $g_{0}$ does not satisfies $43t_{0}\leq g_{0}<8t_{0}-$ 4, then $G\cong K_{\lfloor t\rfloor}\nabla H$, where $H$ consists of stars as follows:
(i) one maximal star $K_{1,\lfloor 1/t_{0}\rfloor+a}$, where $a=\min\big{\{}\big{\lfloor}\frac{\beta_{0}}{t_{0}-\phi_{0}}\big{\rfloor},\alpha_{0}-1\big{\}}$;
(ii) $b$ big stars $K_{1,\lfloor 1/t_{0}\rfloor}$, where $b=\alpha_{0}-a-1$;
(iii) $\ell$ small stars $K_{1,\lfloor 1/t_{0}\rfloor-1}$, where $\ell=n_{0}-b-1$;
(iv) one minimal star $K_{1,\gamma-1}$, where $\gamma=\gamma_{0}$.
Note that when $a=0$, it indicates that there is no maximal star in $G[U]$. Consequently, the number of big stars in $G[U]$ simplifies to $\alpha_{0}$. In the case $\gamma_{0}=\lfloor 1/t_{0}\rfloor$, it implies that there is no minimal star in $G[U]$. The absence of a minimal star changes the composition of small stars. Specifically, the number of small stars becomes $n_{0}-b$.
In summary, by using Theorems 4.1 and 4.2, we have successfully completed the proof of Theorem 1.3. To conclude this section, we will present a corollary.
If $t$ is a positive integer, then $t_{0}=\lfloor t\rfloor+1-t=1$ and $\phi_{0}=(1+\lfloor 1/t_{0}\rfloor)t_{0}-1=1$. Based on (19), we further have $\alpha_{0}=\lfloor g_{0}\rfloor.$ It follows that $a=\alpha_{0}-1=\lfloor g_{0}\rfloor-1$ and $b=0$. By applying Theorems 4.1 and 4.2, we obtain the following corollary.
Corollary 4.3.
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. If $t$ is a positive integer, then $G\cong K_{t}\nabla\big{(}K_{3}\cup(n-t-3)K_{1}\big{)}$ for $43\leq g_{0}<$ 4, or $G\cong K_{t}\nabla\big{(}K_{1,\lfloor g_{0}\rfloor}\cup(n-t-\lfloor g_{0}\rfloor-1)K_{1}\big{)}$ otherwise.
For $t$ being a positive integer, Guiduli proved the Theorem 1.1, and he also conjectured that the spectral extremal graph with respect to Theorem 1.1 is $K_{t}\nabla F$ where the edges within $F$ form a star. Corollary 4.3 resolves this conjecture, and provides a complete characterization of the extremal graphs.
5 Concluding remarks
We may also consider a generalization of the Hereditarily Bounded Property, where the hereditary bound on the number of edges is not restricted to linear constraints. In his thesis, Guiduli posed the following problem.
Problem 5.1.
Given a graph $G$, if $|E(H)|\leq c\cdot|V(H)|^{2}$ holds for all $H\subseteq G$, does it imply that $\rho(G)\leq 2c\cdot|V(G)|$? What can be said if the exponent of $2$ is replaced by some other constant less than $2$?
This issue has also been highlighted by Liu and Ning in their survey of unsolved problems in spectral graph theory [16]. It is worth noting that $\rho(G)\leq\sqrt{2c}n$ serves as a trivial upper bound, and that $2c$ would be best possible, as seen by Wilf’s inequality [20], i.e., $\rho(G)\leq(1-\frac{1}{\omega})n$, where $\omega$ denotes the clique number of $G$. If this question were true, then the spectral Erdős-Stone-Simonovits theorem [17] would follow from the Erdős-Stone-Simonovits theorem [9, 10] and Wilf’s theorem would follow from Turán’s theorem [19].